### v1

# Towards Imperceptible Backdoor Attack in Self-supervised Learning

Abstract
Self-supervised learning models are vulnerable to backdoor attacks. Existing backdoor attacks that are effective in self-supervised learning often involve noticeable triggers, like colored patches, which are vulnerable to human inspection. In this paper, we propose an imperceptible and effective backdoor attack against self-supervised models. We first find that existing imperceptible triggers designed for supervised learning are not as effective in compromising self-supervised models. We then identify this ineffectiveness is attributed to the overlap in distributions between the backdoor and augmented samples used in self-supervised learning. Building on this insight, we design an attack using optimized triggers that are disentangled to the augmented transformation in the self-supervised learning, while also remaining imperceptible to human vision. Experiments on five datasets and seven SSL algorithms demonstrate our attack is highly effective and stealthy. It also has strong resistance to existing backdoor defenses. Our code can be found at https://github.com/Zhang-Henry/IMPERATIVE.
1 Introduction
In recent years, Self-Supervised Learning (SSL) has emerged as a powerful paradigm in machine learning, particularly within the computer vision domain. The key advantage of SSL lies in its ability to learn abundant representations from the vast amounts of unlabeled data, bypassing the labor-intensive and costly process of manual labeling. The objective of SSL in computer vision is to develop an image encoder, designed to ensure that similar images yield close embeddings in the encoder. This is achieved by constructing similar images by adding different augmentations on the same image. Utilizing this pre-trained encoder, users can conduct different downstream tasks by training downstream classifiers based on it. The downstream classifiers are typically compact networks with a small number of parameters.
Although SSL has been extensively used in the development of foundational models [5, 22, 4], it is at risk of backdoor attacks [29, 61, 37, 75], where the attacker embeds hidden malicious behavior within the encoder. The backdoor can be inherited to the downstream task. That is, the downstream classifier will predict a specific target label if the input contains a pre-defined backdoor trigger. Existing backdoor attacks on SSL such as BadEncoder [30] achieve high attack success rates. However, a common drawback of these effective attacks is that their trigger patterns are obvious, making them susceptible to human inspection. In this paper, we aim to propose a backdoor attack in SSL that is both effective and stealthy to human vision.
There are various invisible triggers designed for the backdoor attacks on supervised classifiers, such as WaNet [51], ISSBA [41], and filter attack [46]. A straightforward way to achieve imperceptible backdoor attacks in SSL is by directly applying these invisible triggers. However, these existing invisible triggers designed for supervised learning do not perform as well in attacking self-supervised models (Please refer to section 3 and Figure 3 for more details). We then find that this lack of effectiveness is due to the overlapping distributions between the backdoor samples and the augmented samples utilized in SSL. Namely, self-supervised models cannot effectively distinguish the distribution of the backdoor samples and the augmented samples, due to the similarity between the transformation altered by the backdoor trigger and intrinsic image augmentations in SSL, such as RandomGrayscale and ColorJitter (Please refer to section 3 and Figure 3).
Figure 1: Comparison of clean, backdoored samples created by Patch trigger [18], Instagram filter trigger [33], ISSBA trigger [41], WaNet trigger [51] and ours. Their attack success rates tested under the threat model of BadEncoder [30] are beside the name of the methods. Residuals are the difference between clean and backdoored images.
Based on the above observations, we developed a backdoor attack that disentangles its optimized trigger transformation and the augmented transformation in self-supervised learning. In detail, it involves increasing the distributional distance between backdoor samples and the augmented samples in the self-supervised learning process. we also keep the stealthiness of the trigger by adding the constraints on both pixel-space and feature-space distance to the original samples, using metrics like SSIM, PSNR, perceptual loss, and Wasserstein distance. We then implemented our prototype IMPERATIVE (Imperceptible BAckdoor ATtack In Self-superVised LEarning) and tested it on five datasets (i.e., CIFAR10 [34], STL10 [9], GTSRB [68], SVHN [50], ImageNet [60]), and across seven classic SSL frameworks (i.e., SimCLR [5], MoCo [22], BYOL [17], SimSiam [6], DINO [4], SwAV [3] and CLIP [58], please refer to section 5 and Appendix A) with their various augmentation transforms. The results demonstrate that our method is highly effective and stealthy. In detail, it achieves 99.54% average ASR and 0.95 SSIM. As shown in Figure 1, we compare the backdoor residuals and attack success rates of several methods. Our method exhibits the highest attack success rate while maintaining the highest stealthiness. It also effectively bypasses existing backdoor defenses such as DECREE [15], Beatrix [48], STRIP [16], Grad-CAM [64], Neural Cleanse [83], and commonly studied noises such as JPEG compression, Poisson noise, and Salt&Pepper noise.
Our contributions are summarized as follows: ① We observed that existing imperceptible triggers designed for supervised classifiers have limited effectiveness in SSL. ② We find that the reason behind such ineffectiveness is the coupling feature-space distributions for the backdoor samples and augmented samples in the SSL models. ③ Based on our findings, we propose an imperceptible and effective backdoor attack in SSL by disentangling the distribution of backdoor samples and augmented samples in SSL, while constraining the stealthiness of the triggers during the optimization process. ④ Extensive experiments on five datasets and seven SSL algorithms with their different augmentation ways demonstrate our attack is effective and stealthy, and can also be resilient to current SOTA backdoor defense methods.
2 Related Work
2.1 Self-Supervised Learning
The objective of SSL is to utilize a vast array of unlabeled data (referred to as the pre-training dataset) to pre-train an image encoder. After the pre-training phase, this encoder can then be used to develop classifiers for various downstream tasks utilizing a smaller set of labeled data (known as the downstream dataset) [71]. SSL typically includes the following pipelines for contrastive learning as outlined in [27, 19]: ① Negative Examples [5, 22, 98, 77, 78]: This approach fosters closeness among positive examples and seeks to maximize the distance between negative examples within the latent space. Notable methods include SimCLRs [5] and MoCo [22]. ② Self-distillation [4, 6, 17, 78]: This strategy involves the use of two identical deep neural networks, termed siamese networks, which share the same architecture but operate with distinct weights. The aim is to enhance the similarity between two differently augmented versions of the same image under certain conditions, such as BYOL [17], SimSiam [6], and DINO [4]. ③ Clustering [3]: This approach applies a clustering mechanism by employing swapped prediction of the representations derived from both encoders in an end-to-end framework, such as SwAV [3]. We will demonstrate that our method is highly effective and stealthy under the SSL algorithms in each category in section 5 and Appendix A.
2.2 Backdoor Attacks
Numerous strategies for backdoor attacks in SL (supervised-learning) have been introduced [36, 88, 49, 74, 56, 95, 40], which can be classified based on (i) the objectives of the attacks, including attacks tailored to specific inputs [65, 41], specific classes [69], or all inputs [18]; (ii) the methods of attack, such as poisoning the training data [45], identifying susceptible architectures [54], or distributing models with embedded backdoors [28]; and (iii) the criteria for optimization, focusing on the attack’s effectiveness [53], its ability to be transferred [93], or its stealthiness [80, 7, 99, 65]. Since many attacks mentioned above cannot be applied to SSL due to the necessity of labeled data, recent studies have investigated alternative methods for implanting backdoors into models trained with SSL [61, 2, 43, 30]. However, they usually use a patch as a backdoor trigger and are susceptible to human inspection and model simulation.
2.3 Backdoor Defenses
Various defenses have been developed to defend against backdoor attacks [35, 90, 24, 47, 87, 57, 55, 52, 94, 70]. Most of them are designed to protect against backdoor threats targeting supervised classifiers. These mechanisms work by either preventing attacks during the model training process [85, 25, 23, 79], or by identifying and addressing backdoors in already compromised models in an offline manner [73, 100, 72, 86, 92, 66, 46, 39]. Additionally, some methods are aimed at recognizing backdoor-triggered inputs at the time of inference [8, 11, 16]. DECREE [15] and Beatrix [48] are two strategies specifically designed for SSL, and we will demonstrate that our method can bypass them in subsection 5.4.
3 Observations and Insights
Figure 2: Existing imperceptible backdoor triggers, which yield high ASR in supervised learning (SL), do not perform as effectively in self-supervised learning (SSL). The attack framework for SL and SSL are standard backdoor poisoning [18] and BadEncoder [29], respectively.
Figure 3: SimCLR [5] pre-trained model struggled to differentiate between the backdoor samples with WaNet trigger [51] and augmented samples in SimCLR contrastive learning framework.
We first evaluate the effectiveness of existing invisible triggers (i.e., WaNet [51], ISSBA [41], and filter attack [46]) designed for supervised classifiers. In Figure 3, we show their attack success rates on supervised classifiers and self-supervised models. For the experiments on supervised classifiers, the attack method is standard backdoor poisoning [18]. For self-supervised models, we use the attack method introduced in BadEncoder [30] and replace the patch trigger with these invisible triggers. The model used here is ResNet18 [21]. We observed that these invisible triggers, which yield high Attack Success Rates (ASR) in supervised learning, do not perform as effectively in self-supervised learning. In detail, we employed a ResNet18 pre-trained with SimCLR [5] to train and test a binary classification task, aiming to discern whether samples poisoned by WaNet [51] and those augmented through SimCLR’s contrastive learning-based data augmentation are distinguishable. Throughout this process, we maintained consistent feature representations for clean samples between the backdoored and clean models by using utility loss introduced in Jia et al. [32].
Results indicate that the models struggled to differentiate between the two categories. A t-SNE [81] visualization of their features, as presented in Figure 3, indicates a significant overlap and entanglement. From this, we infer that the diminished effectiveness of supervised backdoor attack methods in self-supervised learning scenarios may be attributed to the distributional similarity between the features of contrastive-learning-augmented samples and backdoor samples. Consequently, our method endeavors to disentangle the feature space distribution of these two sets. The experimental results demonstrate that this approach substantially improves ASR in self-supervised learning.
4 Attack Design
4.1 Threat Model
Objectives of the Attacker.  The objective of an attacker is to implant backdoors into a pre-trained image encoder by SSL (self-supervised learning). We define a backdoored image encoder model as $\mathcal{F}_{\theta}$ and the backdoor injector as $\mathcal{I}_{\phi}$. In this way, a downstream classifier trained based on $\mathcal{F}_{\theta}$, which we define as $\mathcal{C}_{\epsilon}$, could produce a specific prediction $c$ designated by the attacker for inputs $\boldsymbol{x}$ implanted with a trigger chosen by the attacker. The formal definition is shown as follows. $y$ here means the correct label of the input $\boldsymbol{x}$.
$\mathcal{C}_{\epsilon}(\mathcal{F}_{\theta}(\boldsymbol{x}))=y,\quad\mathcal{C}_{\epsilon}(\mathcal{F}_{\theta}(\mathcal{I}_{\phi}(\boldsymbol{x})))=c$ (1)
The ultimate aim of the attacker is to modify a clean image encoder and create a backdoored version that accomplishes several key objectives: ① Effectiveness: The backdoored model should achieve a high rate of successful attacks while preserving high accuracy in benign conditions. ② Naturalness and Stealthiness: To avoid detection during human inspection, the sample with the trigger should appear authentic and natural. ③ Robustness: The effectiveness of the attack should remain intact even when the triggered samples undergo common pre-processing steps, such as samples under various noises. Furthermore, the attack should bypass prevailing defense mechanisms.
Attacker’s Knowledge and Capabilities.  We assume that the attacker has access to a clean image encoder pre-trained following [29]. Additionally, it is presumed that the attacker has access to a collection of unlabeled images, referred to as shadow dataset. The attacker is also assumed to have access to a few images from the Internet, called reference inputs, for each combination of a target downstream task and a target class. We assume that the attacker is capable of manipulating the training procedure to create an encoder with embedded backdoors. Accordingly, the attacker also has access to the augmentation transforms used to pre-train the encoder, which can be utilized in IMPERATIVE to generate stealthy and effective backdoor triggers. However, we assume that the attacker cannot interfere with the training process of these downstream classifiers, such as the training dataset, model framework, and weights.
4.2 Preliminaries
HSV and HSL Color Spaces.  HSV and HSL color spaces align more closely with human color perception, unlike RGB’s additive mixing. They allow for intuitive adjustments of tint, shade, and tone through single-dimension modifications. They have wide applications in the computer vision domain [62, 76, 42, 13]. More details are shown in subsection D.5.
Wasserstein Distance.  Wasserstein distance is commonly employed for quantifying differences between two latent distributions without known common support or density functions [71]. Specifically, we focus on the 2-Wasserstein distance, which is elaborated in subsection D.5.
4.3 Overarching Idea
According to the observations in section 3, the key to enhancing the attack success rate in SSL is to disentangle the two overlapping distributions of the backdoor and augmentation transformation in the contrastive learning of the SSL’s pre-training stage. To widen the distributional gap, we initially convert the image from RGB color space to HSV and HSL color spaces, because they align better with the characteristics of augmentative transformations, such as “ColorJitter” (a data augmentation technique that randomly alters the brightness, contrast, saturation, and hue of an image). Since we already know the augmentation ways in the pre-training stage, thus we design $\mathcal{L}_{\text{disentangle}}$ to quantify the distributional gap between images in the two batches, which involves measuring the difference in color characteristics. To further enlarge the distributional gap, we design $\mathcal{L}_{\text{alignment}}$ to pull close the features of backdoor images and reference images.
Moreover, while we try to expand the distributional difference between the backdoor trigger and the augmentation transformation, an excessively large gap might result in a significant divergence between the backdoored image and the original one. This could, in turn, diminish the naturalness of the backdoored image and reduce the stealthiness of the trigger. Hence we design $\mathcal{L}_{\text{stealthy}}$ to blend the backdoor seamlessly with the original image.
4.4 Our Approach: IMPERATIVE
In our context, we refer to a clean pre-trained image encoder and its backdoor-injected one as $\mathcal{F}_{\theta}$ and $\mathcal{F}_{\theta}^{\prime}$. Given any pair of a downstream task and its corresponding target class, labeled as $\left(T_{i},y_{i}\right)$, the attacker gathers a collection of reference inputs denoted by $R_{i}=\{x_{i1},x_{i2},\ldots,x_{ir_{i}}\}$ from the specified target class $y_{i}$, where $r_{i}$ represents the number of reference inputs for $(T_{i},y_{i})$, with $i=1,2,\cdots,t$. Moreover, for each pair $\left(T_{i},y_{i}\right)$, the attacker chooses a trigger $\boldsymbol{e}_{i}$ to implant into samples in the shadow dataset $\mathcal{D}_{s}$. We denote a clean input $\boldsymbol{x}$ embedded with a trigger as $\boldsymbol{x}^{\prime}$, which is called a backdoored input.
4.4.1 Enhancing Distributional Gap Between Backdoor Images and Augmented Images
To effectively enlarge the distributional distinction between the backdoor and the augmentation in SSL, we design $\mathcal{L}_{\text{disentangle}}$ for scenarios where augmentation transformations might weaken or obscure the pattern of backdoor triggers, leading to a decrease in the ASR. Utilizing $\mathcal{L}_{\text{disentangle}}$ ensures that the distinctiveness of the backdoor is maintained even in the face of various image transformations. The disentangle loss is defined as follows.
$\begin{array}[]{rl}\mathcal{L}_{\text{disentangle}}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum\limits_{\boldsymbol{x}\in D_{s}}&\|H\left(\boldsymbol{x}^{\prime}\right)-H(\tilde{\boldsymbol{x}})\|_{2}+\|S\left(\boldsymbol{x}^{\prime}\right)-S(\tilde{\boldsymbol{x}})\|_{2}\\ +&\|V\left(\boldsymbol{x}^{\prime}\right)-V(\tilde{\boldsymbol{x}})\|_{2}+\|L\left(\boldsymbol{x}^{\prime}\right)-L(\tilde{\boldsymbol{x}})\|_{2},\end{array}$ (2)
where $H$,$S$,$V$,$L$ denote Hue, Saturation, Value, and Lightness transformations from HSV and HSL color spaces, which are elaborated on in subsection 4.2. We denote an input $\boldsymbol{x}$ augmented by the transformations used in the encoder’s pre-training stage as $\tilde{\boldsymbol{x}}$. $\left|\mathcal{D}_{s}\right|$ denotes the sample number in the shadow dataset. $\|u-v\|_{2}$ denotes the $\ell_{2}$ distance between sample $u$ and sample $v$.
4.4.2 Feature Alignment Between Backdoored and Reference Images
Follwoing BadEncoder [30], we enhance the backdoor attack effectiveness by making the compromised image encoder outputs similar feature embeddings for any sample injected with backdoor $\boldsymbol{x}^{\prime}$ in the shadow dataset $\mathcal{D}_{s}$ and the reference inputs $\mathcal{R}_{i}$ of a pair $\left(T_{i},y_{i}\right)$. Consequently, a compromised downstream classifier developed from our compromised image encoder is inclined to assign identical labels to both reference samples $\mathcal{R}_{i}$ and to any compromised sample $\boldsymbol{x}^{\prime}$. We call this process feature alignment between backdoored and reference images, and the $\mathcal{L}_{alignment}$ is defined as follows.
$\mathcal{L}_{alignment}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}^{\prime}\right),\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right)\right)}{\left|\mathcal{D}_{s}\right|\cdot\sum_{i=1}^{t}r_{i}},$ (3)
where $s(\cdot,\cdot)$ is used to quantify the degree of similarity, for instance, cosine similarity, between a pair of feature embeddings. The term $|D_{s}|$ denotes the count of samples within the shadow dataset, and the denominators serve the purpose of standardizing the losses.
4.4.3 Preserving Covert and Natural Backdoors with Expanded Distributional Gaps
We employ several metrics that measure the similarity between the backdoored image and the original one in both pixel and feature space to ensure that our trigger remains both natural and inconspicuous. For assessing similarity in pixel space, we use SSIM [89] and PSNR [26]. Meanwhile, for high-level feature space comparisons, we first use LPIPS [96] which better reflects the subjective experience of image quality and similarity, capturing nuances that simpler metrics might miss. Moreover, since our objective is to reduce the distributional disparity between backdoored and clean samples, the 2-Wasserstein distance (WD) [82] serves as an effective tool for this purpose. We first extract features from backdoored and clean samples respectively using a pre-trained model $\mathcal{M}$, such as ResNet34 [21]. Then we use WD to measure the distance between the extracted features. $\mathcal{L}_{\text{stealthy }}$ is defined as follows.
$\begin{array}[]{rl}\mathcal{L}_{\text{stealthy }}=\sum_{\boldsymbol{x}\in D_{s}}&\lambda_{1}\cdot\mathrm{LPIPS}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right)+\mathrm{WD}\left(\mathcal{M}(\boldsymbol{x}^{\prime}),\mathcal{M}(\boldsymbol{x})\right)\\ -&\lambda_{2}\cdot\mathrm{PSNR}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right)-\mathrm{SSIM}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right),\end{array}$ (4)
where $\lambda_{1},\lambda_{2}$ are used to scale different loss terms to the same scale from 0 to 1.
Algorithm 1 Our backdoor attack IMPERATIVE
Input: Pre-trained benign encoder $\mathcal{F}_{\theta}^{*}$, shadow dataset $\mathcal{D}_{s}$, reference input set $\mathcal{R}$
      Output: Backdoored encoder $\mathcal{F}_{\theta}$ , backdoor trigger injector $\mathcal{I}_{\phi}$
1:function Ours($\mathcal{F}_{\theta}^{*}$, $\mathcal{D}_{s}$ , $\mathcal{R}$)
2:     $\mathcal{F}_{\theta}\leftarrow\mathcal{F}_{\theta}^{*}$; $\hat{\mathcal{D}_{s}}\leftarrow$ Augment samples in $\mathcal{D}_{s}$;
3:     ϕsubscriptitalic-$\mathcal{I}_{\phi}\leftarrow$ a pre-trained backdoor injector $\mathcal{I}_{\phi}^{*}$ using Algorithm 2
4:     for $\text{iter}=0$ to max_epochs do
5:         ${\mathcal{D}_{s}}^{{}^{\prime}}\leftarrow\mathcal{I}_{\phi}(\mathcal{D}_{s})$
6:         $\mathcal{L}_{\text{disentangle}}\leftarrow$ distribution difference between backdoor images $\boldsymbol{x}^{\prime}$ and augmented images $\tilde{\boldsymbol{x}}$, $\forall\boldsymbol{x}^{\prime}\in{\mathcal{D}_{s}}^{{}^{\prime}},\forall\tilde{\boldsymbol{x}}\in\hat{\mathcal{D}_{s}}$triangleright Equation 2
7:         $\mathcal{L}_{\text{stealthy}}\leftarrow$ distance between backdoor image $\boldsymbol{x}^{\prime}$ and clean image $\boldsymbol{x}$, $\forall\boldsymbol{x}^{\prime}\in{\mathcal{D}_{s}}^{{}^{\prime}},\forall\boldsymbol{x}\in\mathcal{D}_{s}$
8:         $\triangleright$ Equation 4
9:         $\mathcal{L}_{alignment}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}\left(\boldsymbol{x}^{\prime}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\left|\mathcal{D}_{s}\right|\cdot\sum_{i=1}^{t}r_{i}}$triangleright Equation 3
10:         $\mathcal{L}_{\text{injector}}=\mathcal{L}_{\text{stealthy}}+\alpha\cdot\mathcal{L}_{\text{disentangle}}+\beta\cdot\mathcal{L}_{\text{alignment}}$triangleright Equation 5
11:         $\phi_{\mathcal{I}}=\phi_{\mathcal{I}}-lr_{1}\cdot\frac{\partial\mathcal{L}_{\text{injector}}}{\partial\phi_{\mathcal{I}}}$
12:         $\mathcal{L}_{\text{consistency }}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\sum_{i=1}^{t}r_{i}}$triangleright Equation 6
13:         $\mathcal{L}_{\text{utility}}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}(\boldsymbol{x}),\mathcal{F}_{\theta}(\boldsymbol{x})\right)$$\triangleright$ Equation 7
14:         $\mathcal{L}_{\text{encoder}}=\mathcal{L}_{\text{alignment}}+\mathcal{L}_{\text{consistency}}+\mathcal{L}_{\text{utility}}$
15:         $\theta_{\mathcal{F}}=\theta_{\mathcal{F}}-lr_{2}\cdot\frac{\partial\mathcal{L}_{\text{encoder}}}{\partial\theta_{\mathcal{F}}}$
4.4.4 Optimization Problem Formulation and Algorithm
We have already defined $\mathcal{L}_{\text{disentangle}}$, $\mathcal{L}_{\text{stealthy }}$, $\mathcal{L}_{alignment}$ in the sections above. Then we can define our IMPERATIVE as an optimization problem. Concretely, our backdoor trigger injector $\mathcal{I}_{\phi}$ is a solution to the subsequent optimization problem:
$\min_{\mathcal{I}_{\phi}}\mathcal{L}=\mathcal{L}_{\text{stealthy }}+\alpha\cdot\mathcal{L}_{\text{disentangle}}+\beta\cdot\mathcal{L}_{\text{alignment}},$ (5)
where $\alpha$ and $\beta$ serve as hyper-parameters to provide equilibrium among these three loss components. We adopt Algorithm 1 to solve the optimization problem, where we alternatively optimize the backdoor injector and the compromised image encoder, and output the final backdoored encoder $\mathcal{F}_{\theta}$ and backdoor trigger injector $\mathcal{I}_{\phi}$. Additionally, to speed up the optimization process and promote the backdoor attack efficacy, we adopt Algorithm 2 to pre-train a backdoor injector to initialize the injector in Algorithm 1 (Please refer to Appendix D.2). We further conduct ablation study to prove the function of the injector pre-training in Appendix B. Finally, we prove that the performance of our model is not sensitive to the hyperparameters in the Appendix C.2.
4.4.5 Highlight of Our Technical Contributions
We wish to highlight our distinct technical contributions: ①Identifying the reason behind existing triggers’ ineffectiveness: We discover that existing backdoor attacks on SSL, such as BadEncoder [30], attain high ASRs but utilize visible triggers. Moreover, traditional imperceptible triggers designed for supervised classifiers, e.g., WaNet [51], ISSBA [41], do not perform effectively in SSL, owing to the overlap in feature-space distributions between the backdoor samples and the augmented samples used in SSL models. ② Proposing disentanglement strategy to generate stealthy and effective backdoor triggers: Based on our insights, we propose a novel backdoor attack in SSL that effectively maintains the potency, stealthiness, and robustness of the generated trigger by disentangling the distribution of backdoor samples from augmented samples within SSL.
5 Evaluation
In this section, we first evaluate the effectiveness and stealthiness of IMPERATIVE using four datasets in subsection 5.2 and subsection 5.3. Then we evaluate the robustness of our method against various SOTA backdoor defenses and noises of different depths in subsection 5.4 and Appendix C.1. After that, we prove our method’s efficacy in large-scale dataset in subsection 5.5. Furthermore, to prove its generalization capability, we conduct more attacking cases on various representative SSL algorithms and a multi-modal model with various augmentations in Appendix A. Moreover, we conduct the ablation study in Appendix B to prove each component’s function. Finally, we evaluate the sensitivity of parameters in Appendix C.2 and further analyze the performance of our method in Appendix C.3.
5.1 Experimental Setup
Datasets.  We utilize the following four image datasets, i.e. CIFAR10 [34], STL10 [9], GTSRB [68], SVHN [50] to evaluate our method, which are also frequently utilized in research related to backdoor attacks [30, 11, 16, 51]. We elaborate on the details of the datasets in Appendix D.6.1.
Pre-training Dataset Downstream Dataset WaNet CTRL Ins-kelvin Ins-xpro2 Ours
[…]
STL10 CIFAR10 86.77 84.43 10.28 87.19 8.72 86.75 18.63 86.85 16.83 87.11 99.58
GTSRB 76.12 74.45 5.23 77.57 8.17 76.49 72.95 76.71 14.02 75.82 97.97
SVHN 55.35 58.29 16.83 54.29 3.32 56.67 38.03 58.42 18.68 58.62 99.76
CIFAR10 STL10 76.14 72.73 9.78 75.73 16.85 74.89 1.16 74.11 5.91 74.48 95.00
GTSRB 81.84 75.85 5.46 79.94 97.95 78.56 2.50 75.08 42.40 79.15 98.73
SVHN 61.52 54.79 17.99 66.33 40.91 68.49 22.13 68.95 30.91 63.67 98.79
Table 1: Results of CA(%), BA(%), and ASR(%) in comparison experiments. Our IMPERATIVE achieves high ASRs while maintaining the accuracy of the downstream classifiers.
Evaluation Metrics.  To assess the effectiveness of our IMPERATIVE, we employ three metrics following existing works [97, 30]: Clean Accuracy (CA), Backdoored Accuracy (BA), Attack Success Rate (ASR). More details are explained in Appendix D.6.2. To evaluate the stealthiness and naturalness of our IMPERATIVE, we employ three metrics following existing works [31, 84]: SSIM [89], PSNR [26] and LPIPS [96]. In experiments, they measure the similarity between backdoored images and original clean images. Higher SSIM and PSNR and lower LIPIPS indicate better stealthiness and naturalness of the generated backdoored images.
Parameter Setting.  We consider a scenario where the attacker chooses a specific downstream task/dataset, a specific target class, and a specific reference input. For the datasets CIFAR10, STL10, GTSRB, and SVHN, the selected target classes are “airplane”, “truck”, “priority sign”, and “digit one”, respectively. CIFAR10 and STL10 are used as the default pre-training dataset. In the pre-training stage, we employ SimCLR [5] by default to train a ResNet18 [21] model, serving as our image encoder. Furthermore, we prove the effectiveness of our method across various representative SSL frameworks in Appendix A, including MoCo [22], BYOL [17], SimSiam [6], DINO [4], SwAV [3], and CLIP [58]. More details are shown in Appendix D.6.3.
Baselines.  We first choose two normal Instagram filters, i.e., kelvin and xpro2, as our baseline triggers, which are gradient-free methods to apply aesthetic enhancements to images, such as adjusting color balance, contrast, and saturation [33, 46]. Moreover, WaNet [51], and CTRL [37], ISSBA [41] are also chosen as our baseline triggers to compare with our method. They are state-of-the-art backdoor triggers and outperform earlier backdoor triggers in terms of stealthiness, yet maintain a high ASR. They are implemented using the default hyperparameters as specified in their original papers. Note that we use BadEncoder [30] to inject these baseline triggers into the compromised encoders.
5.2 Effectiveness Evaluation
Effective Attack.Table 1 indicates our method consistently achieves high ASRs (most above 98%) across various datasets. For instance, with STL10 as the pre-training dataset and CIFAR10 as the downstream dataset, our method reaches an ASR of 99.58%, indicating a highly successful rate of triggering the desired misclassification. This is significantly higher than the ASRs of other benchmark methods like WaNet [51] and CTRL [37] in the same context.
Accuracy Preservation.Table 1 proves that IMPERATIVE preserves the accuracy of the downstream classifiers. The differences between BA and CA are within 2% across the different combinations of pre-training and downstream datasets. This suggests that the presence of the backdoor introduced by our method does not compromise the classifier’s ability to correctly label clean images. For instance, when STL10 is used for pre-training and GTSRB is the downstream dataset, the BA is 76.12% compared to the CA of 75.82%, showing a negligible impact on the classifier’s performance.
[…]
5.3 Stealthiness Evaluation
As shown in Table 3, our method exhibits strong stealthiness advantages with SSIM values consistently close to 1 and most above 0.95, indicating minimal structural changes to the images. PSNR values are notably high, exceeding 40 for some cases, which implies that the noise introduced is hardly detectable. Furthermore, the LPIPS values are exceptionally low, ranging from 0.0001 to 0.02, demonstrating that the perceptual difference between the original and perturbed images is almost negligible. These figures collectively suggest that our method introduces changes that are virtually imperceptible to both human eyes and standard image quality metrics.
5.4 Robustness Evaluation
To assess the resilience of our method against current backdoor defenses, we deploy various state-of-the-art backdoor defense strategies, i.e., DECREE [15], Beatrix [48], Neural Cleanse [83], STRIP [16], Grad-CAM [64] and measure how effectively our method withstands these defenses. Additionally, to further test the robustness of our method, we evaluate its endurance against the following commonly studied noises, i.e., JPEG compression [10, 12, 20, 14], Poisson noise [1, 91], and Salt&Pepper noise [38, 1, 91]. We show the defense results of STRIP, Grad-CAM, and noises in Appendix C.1.
DECREE. [15]  We initially assess whether our method can elude detection by DECREE [15], a method for identifying Trojan attacks in pre-trained encoders. It judges an encoder to be compromised if the proportion of the reversed trigger’s $\mathcal{L}^{1}$ norm relative to the greatest $\mathcal{L}^{1}$ norm within the input space of the model falls below a threshold, i.e., 0.1 $\mathcal{P}\mathcal{L}^{1}$ Norm [15]. As shown in Table 3, $\mathcal{P}\mathcal{L}^{1}$-Norm of each pre-trained and downstream dataset pair are above a default threshold 0.1, and thus this approach does not succeed in identifying backdoored encoders created by IMPERATIVE. This is because the stealthy loss is designed to narrow the distribution gap between backdoored and normal data in both pixel and feature levels. This makes it challenging for DECREE to detect abnormal model internal representation changes when specific behaviors are activated.
Beatrix. [48]  Beatrix [48] identifies poisoned samples by detecting abnormalities in the feature space. We used two pretraining datasets, CIFAR-10 and STL-10, and created backdoored encoders using BadEncoder and IMPERATIVE. By sampling 500 clean inputs and 500 poisoned samples, we applied Beatrix to differentiate them. The findings, as illustrated in Table 5, show that Beatrix effectively recognizes poisoned samples from BadEncoder with over 93% accuracy. However, Beatrix struggles to identify poisoned samples from IMPERATIVE, with a detection accuracy of below 50% on both CIFAR-10 and STL-10, which is like random guessing. We further analyze the reason why Beatrix fails to detect IMPERATIVE in Appendix C.1.1.
[…]
Neural Cleanse. [83]  We then test the resilience of our method against a well-known reverse engineering defense, namely Neural Cleanse (NC) [83]. Specifically, it utilizes the anomaly index to measure the deviation of the reconstructed triggers by their sizes, labeling models with an anomaly index exceeding two as Trojan-infected. Since NC is specifically designed for classifiers and cannot be directly used on image encoders, NC is employed to identify backdoors in a compromised downstream classifier. Moreover, the testing dataset from a targeted downstream dataset serves as the clean dataset. Experimental outcomes, as illustrated in Table 3, indicate that NC is unable to detect the Trojan model created by our approach. While effective in identifying patch-based Trojans [44], this method presumes a uniform trigger pattern at the pixel level across different samples. Our method circumvents this by employing input-dependent triggers, meaning the pixel-level Trojan alterations vary across samples.
5.5 Generalization to Large-scale Dataset
We the generalization ability of our method to large-scale dataset by applying our attack to an image encoder pre-trained on ImageNet [60] released by Google [5]. We show the detailed experimental setup in Appendix D.6.4. Table 5 indicates that our method is highly effective on ImageNet. It achieves high ASR across different datasets, ranging from 96% to 100%, showing its capability to consistently deceive the target system. As shown in Table 3, the high SSIM values ranging from 0.989 to 0.998, high PSNR values ranging from 32.14 to 38.34, and low LPIPS values demonstrate that the perturbations made by IMPERATIVE are almost imperceptible, maintaining the visual quality of the images. Moreover, the BA is close to CA, indicating our attack maintains accuracy for the given downstream task despite the backdoor. Additionally, both our ASR and BA are much higher than those of the benchmark baseline ISSBA [41], proving ours has better performance.
6 Conclusions and Future Work
In this paper, we propose an imperceptible and effective backdoor attack against self-supervised models based on the optimized triggers that is disentangled to the augmented transformation in the self-supervised learning. Based on the evaluation across five different datasets and seven SSL algorithms, our attack is demonstrated to be both highly effective and stealthy. It also effectively bypasses existing backdoor defenses. For future work, it would be beneficial to expand the scope of research to include various other domains of machine learning, such as NLP and audio processing.
Appendix A More Attacking Cases of IMPERATIVE
A.1 Generalization to Other Representative SSL Frameworks
As elaborated on subsection 2.1, SSL frameworks for Contrastive Learning can be normally based on the following categories: negative examples, self-distillation, and clustering. To further evaluate the efficacy of IMPERATIVE, we test our method across the representative SSL algorithms in each category apart from the default SimCLR, i.e., negative examples (MoCo [22]), self-distillation (BYOL [17], SimSiam [6], and DINO [4]), and clustering (SwAV [3]). Table 7 illustrates the detailed parameters for different augmentation transforms used in the SSL algorithms, including Random resized crop, Random horizontal flip, Color jitter, Random gray scale, and Random polarization. The backbone feature extractor used here is ResNet18 [21]. Other experiment settings are the same as those in Appendix D.6.3.
Table 6 shows the evaluation results of applying IMPERATIVE to the encoders pre-trained utilizing various SSL algorithms on CIFAR-10. The SSL algorithms demonstrate high effectiveness in achieving successful backdoor attacks with ASR nearing or reaching perfection in many cases, with an average ASR of 99.54%. Additionally, the algorithms excel in maintaining the stealthiness of these attacks, evidenced by a high average SSIM of 0.95 and a high average PSNR value of 26.89, alongside a low LPIPS value of 0.011. These metrics across the SSL algorithms collectively indicate that while IMPERATIVE are highly effective at manipulating model behaviors, they do so in a way that remains largely undetectable through standard image quality assessments, which aligns with the conclusions drawn in section 5.
[…]
MOCO [22] STL10 72.56 99.75 0.95 26.74 0.0080
GTSRB 68.11 99.98 0.95 26.84 0.0070
SVHN 74.58 99.97 0.96 31.86 0.0030
SimSiam [6] STL10 57.56 100.00 0.95 23.29 0.0230
GTSRB 28.99 99.90 0.96 24.26 0.0190
SVHN 52.24 100.00 0.96 26.04 0.0141
BYOL [17] STL10 39.31 98.33 0.95 26.94 0.0070
GTSRB 11.15 99.31 0.96 32.35 0.0020
SVHN 42.76 99.74 0.96 31.88 0.0024
SwAV [3] STL10 66.28 99.76 0.89 23.43 0.0258
GTSRB 79.39 100.00 0.97 25.42 0.0089
SVHN 76.71 96.94 0.96 25.05 0.0096
DINO [4] STL10 36.06 99.57 0.93 26.98 0.0141
GTSRB 11.49 99.83 0.95 27.11 0.0069
SVHN 37.66 99.99 0.93 25.18 0.0184
Average 50.32 99.54 0.95 26.89 0.0110
Table 6: Evaluations on other SSL algorithms, which are used to pre-train the encoders on CIFAR-10. IMPERATIVE keeps highly effective and stealthy across these algorithms.
Parameter SimCLR/CLIP MOCO SimSiam BYOL SwAV DINO Description
input_size 32/224 32 32 32 / / Size of the input image in pixels.
cj_prob 0.8 0.8 0.8 0.8 0.8 0.8 Probability that color jitter is applied.
cj_strength 1 0.5 1 1 0.5 0.5 Strength of the color jitter. cj_bright, cj_contrast, cj_sat, and cj_hue are multiplied by this value.
cj_bright 0.4 0.4 0.4 0.4 0.8 0.8 How much to jitter brightness.
cj_contrast 0.4 0.4 0.4 0.4 0.8 0.8 How much to jitter contrast.
cj_sat 0.4 0.4 0.4 0.2 0.8 0.4 How much to jitter saturation.
cj_hue 0.1 0.1 0.1 0.1 0.2 0.2 How much to jitter hue.
min_scale / 0.2 0.2 0.08 / / Minimum size of the randomized crop relative to the input_size.
random_gray_scale 0.2 0.2 0.2 0.2 0.2 0.2 Probability of conversion to grayscale.
vf_prob 0 0 0 0 0 0 Probability that vertical flip is applied.
hf_prob 0.5 0.5 0.5 0.5 0.5 0.5 Probability that horizontal flip is applied.
rr_prob 0 0 0 0 0 0 Probability that random rotation is applied.
crop_sizes / / / / [32,32] / Size of the input image in pixels for each crop category.
crop_counts / / / / [2,2] / Number of crops for each crop category.
crop_min_scales / / / / [0.14,0.14] / Min scales for each crop category.
crop_max_scales / / / / [32,32] / Max_scales for each crop category.
global_crop_size / / / / / 32 Crop size of the global views.
global_crop_scale / / / / / [0.4, 1.0] Tuple of min and max scales relative to global_crop_size.
local_crop_size / / / / / 96 Crop size of the local views.
local_crop_scale / / / / / [0.05, 0.4] Tuple of min and max scales relative to local_crop_size.
n_local_views / / / / / 6 Number of generated local views.
solarization_prob / / / / / 0.2 Probability to apply solarization on the second global view.
Table 7: Parameters of Augmentation Transforms in SSL Algorithms.
A.2 Generalization to Multi-modal Models
We utilize our IMPERATIVE to integrate a backdoor into the image encoder of the multi-modal model CLIP [58], which is comprised of both an image and a text encoder, pre-trained on 400 million internet-sourced (image, text) pairs. This backdoored image encoder is then employed to construct a multi-shot downstream classifier.
Due to the unavailability of CLIP’s original pre-training dataset, we utilize the training images from CIFAR-10, resized to 224×224×3, as our shadow dataset. We define $\alpha=0.1$ and $\beta=20$ in Algorithm 1, and $\mu=0.1$ in Algorithm 2 to normalize the loss terms across similar scales. Additionally, Algorithm 1 and Algorithm 2 are configured with a batch size of 16, and learning rates are set at $10^{-6}$ and 0.005, respectively, for a duration of 100 epochs. SVHN serves as the downstream dataset with the target label “one”. We employ the same reference inputs as described in [29]. All other experimental settings remain consistent with those outlined in Appendix D.6.3.
Table 8 presents the attack results of multi-shot classification using CLIP. IMPERATIVE achieves a similar high BA and ASR as BadEncoder [29] does, but it maintains much more stealthiness with high SSIM&PSNR and low LPIPIS values than BadEncoder which uses a white patch as the backdoor trigger.
Methods CA(%) BA(%) ASR(%) SSIM PSNR LPIPIS
Ours 70.6 68.98 99.95 0.98 30.11 0.0179
BadEncoder [29] 70.27 99.99 / / /
Table 8: Attack results on the multi-shot classification of CLIP. IMPERATIVE attains comparable high BA and ASR to those of BadEncoder [29], yet it offers significantly enhanced stealthiness.
Appendix B Ablation Study
As discussed in subsection 4.4, we utilize $\mathcal{L}_{disentangle}$ to enlarge the gap between the backdoored image distribution and the augmented image distribution. We use $\mathcal{L}_{alignment}$ to promote feature alignment between backdoored images and reference images. We use $\mathcal{L}_{stealthy}$ to make the backdoor trigger stealthy and natural. In Algorithm 1, we initialize a backdoor injector using a pre-trained one by Algorithm 2. We use STL10 as the dataset for pre-training encoders. In this section, we demonstrate the effects of each important component in subsection 4.4 by ablating them respectively.
As reported in Table 9, the ASRs of our method decrease across all target downstream datasets when ablating each component respectively. As reported in Table 10, the average SSIM and PSNR drop dramatically and LPIPS increases a lot across all datasets when ablating $\mathcal{L}_{stealthy}$. For example, in the ablation case of the target downstream dataset CIFAR10, the ASR decreases from 99.58% to 67.84%, 10.22%, 75.27% when ablating $\mathcal{L}_{disentangle}$, $\mathcal{L}_{alignment}$ and backdoor injector initialization respectively. This distinctly shows that the absence of key components significantly compromises the effectiveness of our backdoor attack methodology. Moreover, when ablating $\mathcal{L}_{stealthy}$, the SSIM drops from 0.99 to 0.1. The PSNR drops from 41.80 to 5.65. The LPIPS increases from 0.0002 to 0.55. This indicates that the backdoored images without the $\mathcal{L}_{stealthy}$ are more easily detectable and differ a lot from the original images.
Pre-training Backdoor injector $\mathcal{L}_{\text{disentangle}}$ $\mathcal{L}_{\text{alignment}}$ $\mathcal{L}_{\text{stealthy }}$ Target Downstream Dataset
CIFAR10 GTSRB SVHN
[…]
84.66 75.27 78.72 73.33 58.38 13.87
85.71 67.84 74.05 52.26 56.41 62.56
87.20 10.22 65.49 7.46 56.01 23.44
87.11 99.58 75.82 97.97 58.62 99.76
Table 9: ASR(%), BA(%) in ablation studies. Ablating the $\mathcal{L}_{\text{alignment}}$ means ablating the one in $\mathcal{L}_{\text{injector}}$ instead of $\mathcal{L}_{encoder}$, and ablating the pre-trained backdoor injector means randomly initializing it.
Method Target Downstream Dataset
CIFAR10 GTSRB SVHN
[…]
w/o $\mathcal{L}_{stealthy}$ 0.10 5.65 0.5500 0.02 5.20 0.4400 0.10 4.85 0.52
Ours 0.99 41.80 0.0002 0.99 43.58 0.0001 0.96 25.65 0.01
Table 10: Results of ablating $\mathcal{L}_{stealthy}$. Images backdoored without the $\mathcal{L}_{stealthy}$ are more readily identifiable and exhibit significant deviations from the original images.
Appendix C More Evaluations
C.1 More Robustness Tests
C.1.1 Beatrix
Further examination of Beatrix’s detection performance reveals that it determines whether a sample is poisoned based on the deviation in the feature matrix within the embedding space [48]. A sample is deemed poisoned if the deviation exceeds a specific threshold. Figure 4 displays the distribution of deviation values. Notably, poisoned samples by BadEncoder show significantly higher deviation than clean samples, enabling Beatrix to achieve high detection accuracy. Conversely, the deviation distributions between clean inputs and poisoned samples by IMPERATIVE overlap considerably, making them difficult to differentiate.
Figure 4: Beatrix’s deviation distribution [48] for clean and backdoored data, where blue bars represent clean samples and orange bars indicate poisoned ones. Poisoned samples from BadEncoder display markedly higher deviations than clean samples, allowing Beatrix to detect them effectively. However, the deviation distributions for clean inputs and poisoned samples from IMPERATIVE significantly overlap, complicating their differentiation.
Figure 5: Resilience to Grad-CAM [64]. The resemblance observed in these heatmaps indicates that IMPERATIVE is capable of resisting defenses that rely on Grad-CAM.
C.1.2 Grad-CAM
We assess IMPERATIVE’s resilience to Grad-CAM [64], which generates a heatmap for a given model and input sample, where the heat intensity of each pixel reflects its significance in the model’s final prediction. Grad-CAM is effective for identifying smaller Trojans [45, 18], as these Trojans tend to produce high heat values in compact trigger zones, leading to abnormal heatmaps. Nevertheless, our backdoor transformation function alters the entire image, rendering Grad-CAM ineffective in detecting it. Figure 5 illustrates the comparison of heatmaps from both clean images and backdoored images created by our method. The similarity in these heatmaps suggests that IMPERATIVE can withstand defenses based on Grad-CAM.
C.1.3 Various Noises
Next, we are going to introduce the noises, i.e., JPEG compression, Salt&Pepper noise, and Poisson noise, applied in the robustness test.
  • JPEG Compression: JPEG is a prevalent image format. It is a common case that images undergo compression, especially during web transmission. The quality scale ranges from 1 to 100, with 100 being the highest quality and 1 being the lowest quality.
  • Poisson Noise: It arises from the statistical nature of photon reception by image sensors, resulting in random noise of varying intensities. It is set between 1 and 10 to model different lighting conditions and sensor sensitivities.
  • Salt&Pepper Noise: This noise model mimics random pixel corruption due to image sensor errors, transmission faults, or system failures, characterized by randomly scattered white (salt) and black (pepper) pixels. It typically ranges from 0 to 1, but practical applications often use a much smaller range (e.g., 0.01 to 0.2) to avoid overwhelming the image with noise.
Figure 6: Resilience to JPEG compression, Salt&Pepper noise, Poisson noise of different densities.
Evaluating the robustness of our method across these noises helps ensure its performance is reliable even when image quality is compromised when encountering real-world image issues. We illustrate the evaluation results in Figure 6, where the pre-trained dataset is CIFAR10 and the downstream dataset is STL10. The results prove that our method, which utilizes $\mathcal{L}_{disentangle}$, maintains high ASRs in most cases, and the ASR is significantly higher than the cases ablating $\mathcal{L}_{disentangle}$ in our method. For example, under different intensities of Poisson noise, the ASR of our method maintains nearly 100%, while the ASR of the cases ablating $\mathcal{L}_{disentangle}$ in our method is unstable, ranging from 12% to 61%. The outcome indicates that the $\mathcal{L}_{disentangle}$ plays a prominent role in promoting the robustness of our method. The reason may be that the $\mathcal{L}_{disentangle}$ encourages the model to learn features that remain invariant under the transformations. Namely, the model can still recognize the embedded triggers even when various noises or other image perturbations are applied. This aspect of feature preservation is key to ensuring that the backdoor attack remains effective regardless of the image transformations or certain noises applied.
C.1.4 STRIP
Figure 7: Resilient to STRIP [16]. IMPERATIVE can withstand the STRIP runtime defense.
STRIP [16] scrutinizes a suspect sample by overlaying diverse image patterns onto it and monitoring the uniformity of the model’s predictions for these altered inputs. A low entropy score, suggesting consistent predictions across these perturbed samples, would lead STRIP to classify the sample as Trojan-infected. The experimental outcomes depicted in Figure 7 reveal that the entropy values for both clean and Trojan-compromised models produced through our method overlap considerably, signifying that our attack can withstand the STRIP runtime defense. IMPERATIVE manages to evade STRIP as the superimposition process destroys the attack’s trigger, thereby causing both the Trojan-induced and clean samples to exhibit significant alterations in prediction when superimposed, aligning them with what is observed in clean cases.
C.2 Sensitivity Evaluation
Figure 8: Sensitivity evaluation of hyperparameters. IMPERATIVE performs consistently over these parameter ranges.
Figure 8 illustrates the sensitivity evaluation of the hyperparameters alpha and beta, showcasing that the model’s performance exhibits relative insensitivity to parameter variations. Specifically, for alpha, both the Attack Success Rate (ASR) and Backdoored Accuracy (BA) remain relatively constant across a range of alpha values from 0.05 to 0.30, suggesting that once an optimal threshold is reached, the model’s performance is stable despite further adjustments to alpha. Similarly, for beta, changes across a broader spectrum, from 0 to 10, show that ASR quickly stabilizes at 100%, while BA and CA display minimal fluctuations, indicating the robust model performance that is not significantly affected by variations in beta. The horizontal axis in both graphs represents the incremental values of alpha and beta, demonstrating the model’s consistent performance over these parameter ranges.
C.3 Results Analysis
Figure 9: Visualization of feature space segregation over epochs demonstrates the efficacy of our backdoor trigger in creating a distinguishable separation between backdoor and clean features, unlike WaNet [51] where the features remain intermingled, highlighting our method’s utility in securing encoders against backdoor threats.
Figure 9 presents a visual analysis of the feature shifts induced by backdoor injection across different epochs for our proposed method compared to the WaNet [51]. Throughout the backdoor injection process, it becomes evident that our method’s trigger effectively separates the backdoor features from the clean features. This distinction is observable as the training progresses from epoch 10 through to epoch 200, where the separation between the backdoor and clean features becomes increasingly pronounced, indicating a clear demarcation that our encoder can exploit to differentiate between the two.
In contrast, the bottom row representing WaNet’s performance shows a significant overlap of backdoor and clean features. Even at epoch 200, the features remain intermixed, demonstrating WaNet’s inability to distinguish effectively between the two, as the clean, backdoor, and reference features are all entangled. The effectiveness of our backdoor trigger is thus underscored. It introduces a discernible feature shift that an encoder can leverage to recognize backdoored inputs, validating the practical utility of our approach in enhancing model security against backdoor attacks.
Appendix D Implementation Details
The following sections provide a comprehensive breakdown of the process of implementing the attack, which is methodically organized into three distinct phases: pre-training the encoder (Appendix D.1), pre-training the backdoor injector (Appendix D.2), injecting the backdoor (Appendix D.3), and training the downstream classifiers (Appendix D.4). Additionally, we explain the HSV&HSL Color Spaces and Wasserstein Distance (Appendix D.5) used in the method in detail. Moreover, we provide more details on the datasets (Appendix D.6.1), evaluation metrics (Appendix D.6.2), and experimental setup details (Appendix D.6.3, Appendix D.6.4).
D.1 Pre-training Image Encoders
The initial step involves pre-training an image encoder using a specific dataset, hereinafter referred to as the pre-training dataset. In alignment with the experimental framework established in BadEncoder [30], we opt for CIFAR10 or STL10 as our pre-training datasets. These choices are informed by their relatively large size and superior data quality. Crucially, our experimentation leverages the pre-trained weights from BadEncoder’s image encoders, which underwent training over 1,000 epochs employing the Adam optimizer at a learning rate of 0.001 [29]. Specifically, for the CIFAR10 dataset, the encoder is pre-trained using only the training images, excluding their labels. Meanwhile, for STL10, the pre-training also incorporates its additional unlabeled images.
D.2 Pre-training Backdoor Injector
To enhance the optimization speed and increase the effectiveness of the backdoor attack, we use the method outlined in Algorithm 2 to pre-train a backdoor injector, which serves as the initial setup for the injector in Algorithm 1. We use the U-Net architecture [59] for the backdoor injector. The detailed architecture is shown in the Table 11.
Algorithm 2 Pre-training backdoor injector
Input: Shadow dataset $\mathcal{D}_{s}$       Output: Pre-trained backdoor injector $\mathcal{I}_{\phi}$
1:function Pre-training injector($\mathcal{D}_{s}$)
2:     $\hat{\mathcal{D}_{s}}\leftarrow$ Augment samples in $\mathcal{D}_{s}$; $\mathcal{I}_{\phi}\leftarrow\text{Random initialization}$
3:     for $\text{iter}=0$ to max_epochs do
4:         ${\mathcal{D}_{s}}^{{}^{\prime}}\leftarrow\mathcal{I}_{\phi}(\mathcal{D}_{s})$
5:         $\mathcal{L}_{\text{disentangle}}\leftarrow$ distribution difference between backdoor images $\boldsymbol{x}^{\prime}$ and augmented images $\tilde{\boldsymbol{x}}$, $\forall\boldsymbol{x}^{\prime}\in{\mathcal{D}_{s}}^{{}^{\prime}},\forall\tilde{\boldsymbol{x}}\in\hat{\mathcal{D}_{s}}$triangleright Equation 2
6:         $\mathcal{L}_{\text{stealthy}}\leftarrow$ distance between backdoor image $\boldsymbol{x}^{\prime}$ and clean image $\boldsymbol{x}$, $\forall\boldsymbol{x}^{\prime}\in{\mathcal{D}_{s}}^{{}^{\prime}},\forall\boldsymbol{x}\in\mathcal{D}_{s}$
7:         $\triangleright$ Equation 4
8:         $\mathcal{L}_{\text{ours}}=\mathcal{L}_{\text{stealthy}}+\mu\cdot\mathcal{L}_{\text{disentangle}}$
9:         $\phi_{\mathcal{I}}=\phi_{\mathcal{I}}-lr\cdot\frac{\partial\mathcal{L}_{\text{ours}}}{\partial\phi_{\mathcal{I}}}$
Module Layer Type Kernel Size Stride Padding Channel I/O Normalization Activation Input
Encoder Maxpool MaxPool2d 2x2 2 0 - - - Input Image
Conv1 Conv2d 3x3 1 1 3/8 BatchNorm2d (8) ReLU Maxpool
Conv2 Conv2d 3x3 1 1 8/16 BatchNorm2d (16) ReLU Conv1
Conv3 Conv2d 3x3 1 1 16/32 BatchNorm2d (32) ReLU Conv2
Conv4 Conv2d 3x3 1 1 32/64 BatchNorm2d (64) ReLU Conv3
Conv5 Conv2d 3x3 1 1 64/128 BatchNorm2d (128) ReLU Conv4
Decoder Up5 Upsample + Conv2d -/3x3 -/1 -/1 128/64 BatchNorm2d (64) ReLU Conv5
Up_conv5 Conv2d 3x3 1 1 128/64 BatchNorm2d (64) ReLU Up5
Up4 Upsample + Conv2d -/3x3 -/1 -/1 64/32 BatchNorm2d (32) ReLU Up_conv5
Up_conv4 Conv2d 3x3 1 1 64/32 BatchNorm2d (32) ReLU Up4
Up3 Upsample + Conv2d -/3x3 -/1 -/1 32/16 BatchNorm2d (16) ReLU Up_conv4
Up_conv3 Conv2d 3x3 1 1 32/16 BatchNorm2d (16) ReLU Up3
Up2 Upsample + Conv2d -/3x3 -/1 -/1 16/8 BatchNorm2d (8) ReLU Up_conv3
Up_conv2 Conv2d 3x3 1 1 16/8 BatchNorm2d (8) ReLU Up2
Conv_1x1 Conv2d 1x1 1 - 8/3 - - Up_conv2
Table 11: The detailed network structure of backdoor injector.
D.3 Backdoor Injection
The subsequent phase introduces a backdoor into the pre-trained encoder. This is achieved through the use of Shadow Dataset, which constitutes a smaller subset of the pre-training dataset. Consistent with preceding studies [29], this shadow dataset comprises 50,000 randomly selected training samples. Additionally, we employ the same reference samples as those used in BadEncoder [29]. The backdoor is injected using a designated algorithm, which leverages a combination of $\mathcal{L}_{consistency}$, $\mathcal{L}_{utility}$, and $\mathcal{L}_{alignment}$. The notations are the same as the main paper.
Consistency Loss. The consistency loss decreases when the feature vectors generated by the backdoored image encoder and the clean image encoder for the reference inputs are more alike.
$\mathcal{L}_{consistency}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\sum_{i=1}^{t}r_{i}}$ (6)
Utility Loss. Utility loss diminishes when the backdoored image encoder and the clean image encoder yield more closely matching feature vectors for a clean input within the shadow dataset.
$\mathcal{L}_{utility}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}(\boldsymbol{x}),\mathcal{F}_{\theta}(\boldsymbol{x})\right)$ (7)
D.4 Training Downstream Classifiers
Finally, using the pre-trained image encoder, we proceed to train classifiers for three separate datasets, hereafter referred to as downstream datasets. Our approach utilizes a two-layer fully connected neural network as the classifier architecture, with the first and second layers consisting of 512 and 256 neurons, respectively. The testing subset of each downstream dataset is utilized for evaluation purposes. The training of the classifier is conducted over 500 epochs using the cross-entropy loss function and the Adam optimizer, with a set learning rate of 0.0001.
D.5 Technical Background
HSV Color Space. The HSV color model describes color based on three primary attributes: hue, saturation, and value. Hue (H) represents the color itself, essentially the aspect of color we typically refer to by names like red or yellow. Saturation (S) indicates the vividness or richness of the color. A higher saturation means a more intense, pure color. Value (V), often referred to as brightness, measures the lightness or darkness of the color [67]. HSV color space can be converted from RGB color space through Equation 8, Equation 9, Equation 10, Equation 11 [63].
HSL Color Space. The HSL color model is defined by three cylindrical coordinates: hue (H), saturation (S), and lightness (L), which together capture the subtleties of color. HSL color space can be converted from RGB color space through Equation 8, Equation 9, Equation 12, Equation 13 [63].
$\mathrm{R}^{\prime}=\mathrm{R}/255;$ (8)
$\mathrm{G}^{\prime}=\mathrm{G}/255;$
$\mathrm{B}^{\prime}=\mathrm{B}/255$
$\mathrm{C}_{\max}=\operatorname{MAX}\left(\mathrm{R}^{\prime},\mathrm{G}^{\prime},\mathrm{B}^{\prime}\right)$
$\mathrm{C}_{\min}=\operatorname{MIN}\left(\mathrm{R}^{\prime},\mathrm{G}^{\prime},\mathrm{B}^{\prime}\right)$
$\Delta=\mathrm{C}_{\max}-\mathrm{C}_{\min}$
$\mathrm{H}=\left\{\begin{array}[]{l}60^{\circ}\times\left(\frac{\mathrm{G}^{\prime}-\mathrm{B}^{\prime}}{\Delta}\bmod 6\right),\mathrm{C}_{\max}=\mathrm{R}^{\prime}\\ 60^{\circ}\times\left(\frac{\mathrm{B}^{\prime}-\mathrm{R}^{\prime}}{\Delta}+2\right),\mathrm{C}_{\max}=\mathrm{G}^{\prime}\\ 60^{\circ}\times\left(\frac{\mathrm{R}^{\prime}-\mathrm{G}^{\prime}}{\Delta}+4\right),\mathrm{C}_{\max}=\mathrm{B}^{\prime}\end{array}\right.$ (9)
$\mathrm{S}=\left\{\begin{array}[]{cc}0\quad,&\Delta=0\\ \frac{\Delta}{\mathrm{C}_{\max}},&\Delta\neq 0\end{array}\right.$ (10)
$\mathrm{V}=\mathrm{C}_{\max}$ (11)
$\mathrm{S}=\left\{\begin{array}[]{cl}0,&\Delta=0\\ \frac{\Delta}{1-|2\mathrm{L}-1|}&,\Delta\neq 0\end{array}\right.$ (12)
$\mathrm{L}=\left(\mathrm{C}_{\max}+\mathrm{C}_{\min}\right)/2$ (13)
Wasserstein Distance.  Wasserstein distance is commonly employed for quantifying differences between two latent distributions without known common support or density functions [71]. Specifically, we focus on the 2-Wasserstein distance, which is defined below.
[…]
where $\xi$ and $\tau$ correspond to the marginal probability distributions of the clean and poisoned datasets. $\psi$ denotes the combined distribution of both clean and poisoned data samples. $\inf$ represents the infimum, which is the lowest boundary value for the calculated distances within the scope of the joint distribution $\psi$. The integral part of the equation aggregates the distances between each pair of data points, $(u,v)$, where one belongs to the clean set and the other to the poisoned set, as drawn from the joint distribution $\psi$. Here, $p(u,v)$ defines the probability of concurrently selecting both data samples. The term $||u-v||_{2}$ measures the $L^{2}$ distance between the two data points [82].
D.6 Experiments Details
D.6.1 Dataset Details
We use the following datasets in our method evaluation.
  • CIFAR10 [34]: This dataset comprises 60,000 images of 32×32×3 pixels and 10 different classes for basic image recognition tasks, divided into 50,000 for training and 10,000 for testing.
  • STL10 [9]: STL10 includes 5,000 labeled training images and 8,000 for testing with a resolution of 96×96×3 pixels, across 10 classes. Additionally, it provides 100,000 unlabeled images for unsupervised learning. Notably, they are resized to 32×32×3 to be consistent with other datasets.
  • GTSRB [68]: This dataset includes 51,800 images of traffic signs categorized into 43 classes. It is split into 39,200 training and 12,600 testing images, each sized at 32×32×3.
  • SVHN [50]: SVHN is a dataset of digit images from house numbers in Google Street View, consisting of 73,257 training and 26,032 testing images, each 32×32×3 in size.
  • ImageNet [60]: ImageNet is designed for large-scale object classification, featuring 1,281,167 training samples and 50,000 testing samples across 1000 categories. Each image has a resolution of 224x224 pixels with three color channels.
D.6.2 Evaluation Metrics Details
The meanings of evaluation metric are shown below:
  • CA: This is the accuracy of a clean downstream classifier in correctly classifying the clean testing images from the corresponding downstream dataset.
  • BA: BA measures the accuracy of a backdoored downstream classifier on the same clean testing images from the corresponding downstream dataset.
  • ASR: ASR measures the percentage of these backdoored test images that are classified as the target class by the backdoored downstream classifier.
D.6.3 Parameter Setting Details
We adopt the default parameter settings as follows unless stated otherwise:
  1. 1.
    We set $\lambda_{1}=10,\lambda_{2}=0.025$ in the loss equation term Equation 4, $\alpha=0.1,\beta=5$ in Algorithm 1, and $\mu=0.1$ in Algorithm 2 to scale the loss terms to the same 0-1 range.
  2. 2.
    The Shadow dataset consists of 50000 images randomly chosen from the pre-training dataset.
  3. 3.
    Cosine similarity is adopted to measure the similarity between two samples’ feature embeddings outputted by an encoder.
  4. 4.
    Since we know the SSL’s pre-training process, the SSL augmentation transforms are also known. By default, the augmentation transforms used are RandomResizeCrop, RandomHorizontalFlip, ColorJitter, and RandomGrayscale.
  5. 5.
    We adopt U-Net [59] as the structure of the backdoor injector, which is widely used in image-to-image tasks.
  6. 6.
    Algorithm 1 and Algorithm 2 utilize the Adam optimizer with 200 training epochs, a batch size of 256, and learning rates of 0.001 and 0.005 respectively.
D.6.4 Experimental Setup in Real World Case ImageNet
We selected a random 1% subset of ImageNet’s training images to serve as the shadow dataset. Additionally, we adjusted the size of each image in both the shadow and downstream datasets to dimensions of 224×224×3 as Google does in the encoder pre-training stage [5]. We set $\alpha=0.1,\beta=20$ in Algorithm 1, and $\mu=0.1$ in Algorithm 2 to scale the loss terms to the same range. Moreover, Algorithm 1 and Algorithm 2 here use a batch size of 16, and learning rates of $10^{-4}$ and 0.005 respectively. Other experiment settings are the same as those in Appendix D.6.3.
Appendix E Ethical Statements
This research is focused on identifying potential weaknesses in SSL that could be exploited by backdoor attacks. Our approach is guided by ethical standards, and we intend to enhance, not compromise, system security. We aim to promote awareness and encourage the creation of strong protective measures by demonstrating current susceptibilities within SSL frameworks.
Appendix F Limitation
The limitation of the current research is its exclusive focus on computer vision (CV) models. This specificity might not address potential vulnerabilities or behaviors applicable to other types of models, such as those used in natural language processing (NLP) or audio signal processing.
Appendix G Broader Impacts
For potential positive societal impacts, first by exposing potential vulnerabilities in SSL models, the research encourages the development of more robust security measures. This can lead to safer deployment of AI systems in critical applications such as healthcare, autonomous vehicles, and financial services, where security is paramount. Moreover, Highlighting the vulnerabilities in AI systems can spur more research into ethical AI practices. This includes developing guidelines and frameworks to ensure that AI technologies are used responsibly and do not inadvertently harm users or perpetuate biases.
For potential negative societal impacts, public knowledge of vulnerabilities in AI systems can lead to a loss of trust in these technologies. If people believe that AI systems can be easily manipulated or compromised, they may be hesitant to rely on these technologies, slowing down technological adoption and potential benefits.

### v2

@@
-Towards Imperceptible Backdoor Attack in Self-supervised Learning
+Invisible Backdoor Attack against Self-supervised Learning
 Abstract
-Self-supervised learning models are vulnerable to backdoor attacks.
-Existing backdoor attacks that are effective in self-supervised learning often involve noticeable triggers, like colored patches, which are vulnerable to human inspection.
-In this paper, we propose an imperceptible and effective backdoor attack against self-supervised models.
-We first find that existing imperceptible triggers designed for supervised learning are not as effective in compromising self-supervised models.
-We then identify this ineffectiveness is attributed to the overlap in distributions between the backdoor and augmented samples used in self-supervised learning.
-Building on this insight, we design an attack using optimized triggers that are disentangled to the augmented transformation in the self-supervised learning, while also remaining imperceptible to human vision.
-Experiments on five datasets and seven SSL algorithms demonstrate our attack is highly effective and stealthy.
+Self-supervised learning (SSL) models are vulnerable to backdoor attacks.
+Existing backdoor attacks that are effective in SSL often involve noticeable triggers, like colored patches or visible noise, which are vulnerable to human inspection.
+This paper proposes an imperceptible and effective backdoor attack against self-supervised models.
+We first find that existing imperceptible triggers designed for supervised learning are less effective in compromising self-supervised models.
+We then identify this ineffectiveness is attributed to the overlap in distributions between the backdoor and augmented samples used in SSL.
+Building on this insight, we design an attack using optimized triggers disentangled with the augmented transformation in the SSL, while remaining imperceptible to human vision.
+Experiments on five datasets and six SSL algorithms demonstrate our attack is highly effective and stealthy.
 It also has strong resistance to existing backdoor defenses.
-Our code can be found at https://github.com/Zhang-Henry/IMPERATIVE.
+Our code can be found at https://github.com/Zhang-Henry/INACTIVE.
 1 Introduction
-In recent years, Self-Supervised Learning (SSL) has emerged as a powerful paradigm in machine learning, particularly within the computer vision domain.
-The key advantage of SSL lies in its ability to learn abundant representations from the vast amounts of unlabeled data, bypassing the labor-intensive and costly process of manual labeling.
-The objective of SSL in computer vision is to develop an image encoder, designed to ensure that similar images yield close embeddings in the encoder.
-This is achieved by constructing similar images by adding different augmentations on the same image.
-Utilizing this pre-trained encoder, users can conduct different downstream tasks by training downstream classifiers based on it.
-The downstream classifiers are typically compact networks with a small number of parameters.
+In recent years, Self-Supervised Learning (SSL) has become a powerful approach in deep learning, enabling the learning of rich representations from vast unlabeled data, thus avoiding manual labeling.
+SSL aims to develop an image encoder that produces similar embeddings for similar images by applying various augmentations to the same image.
+This pre-trained encoder can be used for different downstream tasks by training compact downstream classifiers with relatively few parameters.
 Although SSL has been extensively used in the development of foundational models [6, 20, 5], it is at risk of backdoor attacks [27, 56, 34, 68], where the attacker embeds hidden malicious behavior within the encoder.
 The backdoor can be inherited to the downstream task.
-That is, the downstream classifier will predict a specific target label if the input contains a pre-defined backdoor trigger.
-Existing backdoor attacks on SSL such as BadEncoder [30] achieve high attack success rates.
+The downstream classifier predicts a specific target label if the input contains a pre-defined backdoor trigger.
+Existing backdoor attacks on SSL such as BadEncoder [27] achieve high attack success rates (ASR).
 However, a common drawback of these effective attacks is that their trigger patterns are obvious, making them susceptible to human inspection.
-In this paper, we aim to propose a backdoor attack in SSL that is both effective and stealthy to human vision.
+Moreover, while data-poisoning-based attacks CTRL [34] and BLTO [63] are relatively stealthy, their ASRs are suboptimal, For example, on CIFAR10 CTRL only has 61.90% ASR under BYOL framework and BLTO only has 84.63% ASR under SimSiam framework.
+Furthermore, they also rely on the downstream dataset matching the pre-training dataset distribution, limiting effectiveness across diverse datasets.
+In this paper, we aim to propose a backdoor attack in SSL that is both effective and stealthy to human vision without this distribution dependency.
 There are various invisible triggers designed for the backdoor attacks on supervised classifiers, such as WaNet [47], ISSBA [36], and filter attack [42].
 A straightforward way to achieve imperceptible backdoor attacks in SSL is by directly applying these invisible triggers.
-However, these existing invisible triggers designed for supervised learning do not perform as well in attacking self-supervised models (Please refer to section 3 and Figure 3 for more details).
+However, these existing invisible triggers designed for supervised learning do not perform as well in attacking self-supervised models (see Figure 2).
 We then find that this lack of effectiveness is due to the overlapping distributions between the backdoor samples and the augmented samples utilized in SSL.
-Namely, self-supervised models cannot effectively distinguish the distribution of the backdoor samples and the augmented samples, due to the similarity between the transformation altered by the backdoor trigger and intrinsic image augmentations in SSL, such as RandomGrayscale and ColorJitter (Please refer to section 3 and Figure 3).
-Figure 1: Comparison of clean, backdoored samples created by Patch trigger [18], Instagram filter trigger [33], ISSBA trigger [41], WaNet trigger [51] and ours.
-Their attack success rates tested under the threat model of BadEncoder [30] are beside the name of the methods.
+Namely, self-supervised models cannot effectively distinguish the distribution of the backdoor samples and the augmented samples, due to the similarity between the transformation altered by the backdoor trigger and intrinsic image augmentations in SSL, such as RandomGrayscale and ColorJitter (see Figure 3).
+Figure 1: Comparison of clean, backdoored samples created by Patch trigger used by BadEncoder [27] and DRUPE [68], Instagram filter trigger [30], ISSBA trigger [36], WaNet trigger [47] and ours.
+Except for DRUPE, the ASRs are tested under the threat model of BadEncoder.
 Residuals are the difference between clean and backdoored images.
-Based on the above observations, we developed a backdoor attack that disentangles its optimized trigger transformation and the augmented transformation in self-supervised learning.
-In detail, it involves increasing the distributional distance between backdoor samples and the augmented samples in the self-supervised learning process.
-we also keep the stealthiness of the trigger by adding the constraints on both pixel-space and feature-space distance to the original samples, using metrics like SSIM, PSNR, perceptual loss, and Wasserstein distance.
-We then implemented our prototype IMPERATIVE (Imperceptible BAckdoor ATtack In Self-superVised LEarning) and tested it on five datasets (i.e., CIFAR10 [34], STL10 [9], GTSRB [68], SVHN [50], ImageNet [60]), and across seven classic SSL frameworks (i.e., SimCLR [5], MoCo [22], BYOL [17], SimSiam [6], DINO [4], SwAV [3] and CLIP [58], please refer to section 5 and Appendix A) with their various augmentation transforms.
+Our method achieves the highest ASR while maintaining trigger stealthiness, while other methods either have a much lower ASR or use more easily detectable triggers.
+Figure 2: Existing imperceptible backdoor triggers, which yield high ASR in supervised learning (SL), do not perform as effectively in SSL.
+The attack framework for SL and SSL are standard backdoor poisoning [17] and BadEncoder [27], respectively.
+Based on the above observations, we developed a backdoor attack that disentangles its optimized trigger transformation and the augmented transformation in SSL.
+In detail, it involves increasing the distributional distance between backdoor samples and the augmented samples in the SSL process.
+We also keep the trigger stealthy by adding the constraints on both pixel-space and feature-space distance to the original samples, using metrics like SSIM, PSNR, perceptual loss, and Wasserstein distance.
+We then implemented our prototype INACTIVE (INvisible bACkdoor aTtack In self-superVised lEarning) and tested it on five datasets (i.e., CIFAR10, STL10, GTSRB, SVHN, ImageNet), and across six classic SSL frameworks (i.e., SimCLR [6], MoCo [20], BYOL [16], SimSiam [7], SwAV [4] and CLIP [53] (See Table 1 and Table A1) with their various augmentation transforms (See Table A5).
 The results demonstrate that our method is highly effective and stealthy.
-In detail, it achieves 99.54% average ASR and 0.95 SSIM.
-As shown in Figure 1, we compare the backdoor residuals and attack success rates of several methods.
-Our method exhibits the highest attack success rate while maintaining the highest stealthiness.
-It also effectively bypasses existing backdoor defenses such as DECREE [15], Beatrix [48], STRIP [16], Grad-CAM [64], Neural Cleanse [83], and commonly studied noises such as JPEG compression, Poisson noise, and Salt&Pepper noise.
+In detail, it achieves an average of 99.09% ASR (See Table 1), 0.9763 SSIM, 41.07 PSNR, 0.0046 LIPIS, 0.9751 FSIM, and 13.281 FID (See Table A9).
+As shown in Figure 1, we compare several methods’ backdoor residuals and ASRs.
+Our method exhibits the highest ASR while maintaining the highest stealthiness.
+It also effectively bypasses existing backdoor defenses such as DECREE [13], Beatrix [44], ASSET [48], STRIP [14], Grad-CAM [58], Neural Cleanse [72], and various noise, i.e., JPEG compression, Poisson noise, and Salt&Pepper noise.
 Our contributions are summarized as follows: ① We observed that existing imperceptible triggers designed for supervised classifiers have limited effectiveness in SSL.
@@
 ③ Based on our findings, we propose an imperceptible and effective backdoor attack in SSL by disentangling the distribution of backdoor samples and augmented samples in SSL, while constraining the stealthiness of the triggers during the optimization process.
-④ Extensive experiments on five datasets and seven SSL algorithms with their different augmentation ways demonstrate our attack is effective and stealthy, and can also be resilient to current SOTA backdoor defense methods.
+④ Extensive experiments on five datasets and six SSL algorithms with different augmentation ways demonstrate our attack is effective and stealthy, and can also be resilient to current SOTA backdoor defense methods.
 2 Related Work
 2.1 Self-Supervised Learning
-The objective of SSL is to utilize a vast array of unlabeled data (referred to as the pre-training dataset) to pre-train an image encoder.
-After the pre-training phase, this encoder can then be used to develop classifiers for various downstream tasks utilizing a smaller set of labeled data (known as the downstream dataset) [71].
-SSL typically includes the following pipelines for contrastive learning as outlined in [27, 19]: ① Negative Examples [5, 22, 98, 77, 78]: This approach fosters closeness among positive examples and seeks to maximize the distance between negative examples within the latent space.
-Notable methods include SimCLRs [5] and MoCo [22].
-② Self-distillation [4, 6, 17, 78]: This strategy involves the use of two identical deep neural networks, termed siamese networks, which share the same architecture but operate with distinct weights.
-The aim is to enhance the similarity between two differently augmented versions of the same image under certain conditions, such as BYOL [17], SimSiam [6], and DINO [4].
-③ Clustering [3]: This approach applies a clustering mechanism by employing swapped prediction of the representations derived from both encoders in an end-to-end framework, such as SwAV [3].
-We will demonstrate that our method is highly effective and stealthy under the SSL algorithms in each category in section 5 and Appendix A.
+The goal of SSL is to leverage a large amount of unlabeled data in the pre-training dataset to pre-train an image encoder, which can then be used to create classifiers for various downstream tasks with a smaller set of labeled data [68].
+SSL pipelines for contrastive learning typically include the following approaches [26, 18]: ① Negative Examples: Promotes proximity among positive examples while maximizing the distance between negative examples in the latent space, as seen in SimCLRs [6] and MoCo [20].
+② Self-distillation: Utilizes two identical Siamese networks with different weights to increase the similarity between differently augmented versions of the same image, such as in BYOL [16] and SimSiam [7].
+③ Clustering: Implements a clustering mechanism with swapped prediction of representations from both encoders, as in SwAV [4].
+Our method is shown to be highly effective and stealthy under the SSL algorithms in Table 1 and Table A1.
 2.2 Backdoor Attacks
-Numerous strategies for backdoor attacks in SL (supervised-learning) have been introduced [36, 88, 49, 74, 56, 95, 40], which can be classified based on (i) the objectives of the attacks, including attacks tailored to specific inputs [65, 41], specific classes [69], or all inputs [18]; (ii) the methods of attack, such as poisoning the training data [45], identifying susceptible architectures [54], or distributing models with embedded backdoors [28]; and (iii) the criteria for optimization, focusing on the attack’s effectiveness [53], its ability to be transferred [93], or its stealthiness [80, 7, 99, 65].
-Since many attacks mentioned above cannot be applied to SSL due to the necessity of labeled data, recent studies have investigated alternative methods for implanting backdoors into models trained with SSL [61, 2, 43, 30].
-However, they usually use a patch as a backdoor trigger and are susceptible to human inspection and model simulation.
+Backdoor attacks were initially proposed for supervised learning (SL) to modify a model’s behavior on specific inputs or classes while keeping its general performance intact [33, 77, 45, 67, 85, 51, 84, 15, 2].
+Early backdoor attacks commonly utilized visible triggers like distinctive patches that are easily detectable through visual inspection [17, 59, 64].
+To enhance stealth, subsequent research introduced invisible triggers, which are subtle and blend into the background, helping these attacks evade both human inspection and certain automated defenses [36, 47].
+Since many of these attacks rely on labeled data, recent studies have explored alternative backdoor implantation techniques in SSL models [56, 3, 39, 27, 63].
+However, these typically use a visible backdoor trigger, such as a patch, making them prone to human detection and model simulation.
+The advantage of invisible triggers in SSL is clear: they improve attack stealth, bypassing some conventional defenses that focus on detecting visible anomalies [13, 69].
+However, as we will demonstrate, directly applying existing invisible triggers designed for SL to SSL tasks results in limited attack effectiveness.
+Moreover, backdoor attacks in SSL are generally divided into two types: training-time backdoor injection attacks like BadEncoder [27], which require control of the training in the backdoor injection process, and data-poisoning-based attacks like CTRL [34] and BLTO [63], which rely on poisoned data without needing model specifics.
+BadEncoder modifies a pre-trained encoder to embed triggers that cause targeted misclassifications in downstream tasks.
+By aligning the features of triggered images with an attacker-chosen class, downstream classifiers misclassify triggered inputs while maintaining accuracy on clean data.
+Although CTRL and BLTO are relatively stealthy, they achieve lower ASRs than our method and depend on the downstream dataset matching the pre-training distribution, limiting their versatility across diverse datasets.
+In this paper, we focus on training-time backdoor injection attack due to it has higher attack effectiveness and transferability.
 2.3 Backdoor Defenses
-Various defenses have been developed to defend against backdoor attacks [35, 90, 24, 47, 87, 57, 55, 52, 94, 70].
-Most of them are designed to protect against backdoor threats targeting supervised classifiers.
-These mechanisms work by either preventing attacks during the model training process [85, 25, 23, 79], or by identifying and addressing backdoors in already compromised models in an offline manner [73, 100, 72, 86, 92, 66, 46, 39].
-Additionally, some methods are aimed at recognizing backdoor-triggered inputs at the time of inference [8, 11, 16].
-DECREE [15] and Beatrix [48] are two strategies specifically designed for SSL, and we will demonstrate that our method can bypass them in subsection 5.4.
-3 Observations and Insights
-Figure 2: Existing imperceptible backdoor triggers, which yield high ASR in supervised learning (SL), do not perform as effectively in self-supervised learning (SSL).
-The attack framework for SL and SSL are standard backdoor poisoning [18] and BadEncoder [29], respectively.
-Figure 3: SimCLR [5] pre-trained model struggled to differentiate between the backdoor samples with WaNet trigger [51] and augmented samples in SimCLR contrastive learning framework.
-We first evaluate the effectiveness of existing invisible triggers (i.e., WaNet [51], ISSBA [41], and filter attack [46]) designed for supervised classifiers.
-In Figure 3, we show their attack success rates on supervised classifiers and self-supervised models.
-For the experiments on supervised classifiers, the attack method is standard backdoor poisoning [18].
-For self-supervised models, we use the attack method introduced in BadEncoder [30] and replace the patch trigger with these invisible triggers.
-The model used here is ResNet18 [21].
-We observed that these invisible triggers, which yield high Attack Success Rates (ASR) in supervised learning, do not perform as effectively in self-supervised learning.
-In detail, we employed a ResNet18 pre-trained with SimCLR [5] to train and test a binary classification task, aiming to discern whether samples poisoned by WaNet [51] and those augmented through SimCLR’s contrastive learning-based data augmentation are distinguishable.
-Throughout this process, we maintained consistent feature representations for clean samples between the backdoored and clean models by using utility loss introduced in Jia et al.
-[32].
-Results indicate that the models struggled to differentiate between the two categories.
-A t-SNE [81] visualization of their features, as presented in Figure 3, indicates a significant overlap and entanglement.
-From this, we infer that the diminished effectiveness of supervised backdoor attack methods in self-supervised learning scenarios may be attributed to the distributional similarity between the features of contrastive-learning-augmented samples and backdoor samples.
-Consequently, our method endeavors to disentangle the feature space distribution of these two sets.
-The experimental results demonstrate that this approach substantially improves ASR in self-supervised learning.
+Various defenses have been developed against backdoor attacks [32, 78, 23, 43, 76, 52, 50, 49, 83, 8, 81], primarily targeting supervised classifiers.
+These defenses either prevent attacks during training [74, 24, 22, 70] or detect and mitigate backdoors in compromised models offline [66, 89, 65, 75, 80, 60, 42, 37].
+Some methods also detect backdoor-triggered inputs during inference [14].
+Defense methods like DECREE [13], Beatrix [44], and ASSET [48] are designed for SSL, primarily relying on the visible characteristics of triggers to detect backdoors.
+In contrast, our method uses invisible triggers, effectively bypassing these defenses by breaking their reliance on visual anomalies and making detection more challenging.
+3 Observations and Analysis
+Figure 3: t-SNE visualization of the feature space in the inherent augmentation and backdoor trigger space.
+The SimCLR [6] pre-trained model struggled to differentiate between backdoor samples injected with the WaNet trigger [47] and the augmented samples within the SimCLR contrastive learning framework.
+Invisible Trigger Designed for SL Fails in SSL.
+We first assess the effectiveness of existing invisible triggers (WaNet, ISSBA, and filter attack) designed for supervised classifiers.
+Figure 2 displays their ASRs on both supervised classifiers and self-supervised models.
+For supervised classifiers, the standard backdoor poisoning method [17] is used.
+For self-supervised models, we apply the BadEncoder method, replacing the patch trigger with these invisible triggers, using ResNet18.
+We find that these triggers, which achieve high ASR in supervised learning, are less effective in SSL.
+We then investigate the underlying reason for these results and focus on the following research question: Why does the effective invisible backdoor trigger designed for supervised learning fail on self-supervised learning?
+Cause of the Failure: Entanglement of the Inherent Augmentations and Backdoor Trigger.
+We find that the entanglement of the inherent augmentations in contrastive learning can cause the failure of the backdoor injection with such triggers.
+We provide our analysis in this section.
+One of the core training losses of contrastive learning can be formulated as maximizing the feature space similarity between the augmented samples modified from the same training samples:
+$\arg\max_{\theta_{\mathcal{F}}}s(\mathcal{F}_{\theta}(\mathcal{A}_{1}(\bm{x})),\mathcal{F}_{\theta}(\mathcal{A}_{2}(\bm{x})))$ (1)
+where $s(\cdot,\cdot)$ denotes the similarity measurement, $\mathcal{F}_{\theta}$ is the encoder in training ($\theta_{\mathcal{F}}$ is its parameters), $\bm{x}$ is the training sample, $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are different augmentations sampled from the predefined augmentation space $\mathcal{S}_{\mathcal{A}}$.
+Different from the predefined augmentation space, we also define the learned augmentation space for trained encoders $\mathcal{S}_{\mathcal{A}}^{\prime}$ as the space including a set of transformations where any pair within it can achieve high pairwise similarity on augmented versions of the same sample when processed by the trained encoder.
+We also define perfectly-trained encoder is the encoder that achieves maximal similarity described in Equation 1 for all samples and all possible transformations used, i.e., $s(\mathcal{F}_{\theta}(\mathcal{A}_{1}(\bm{x})),\mathcal{F}_{\theta}(\mathcal{A}_{2}(\bm{x})))=1,\forall\bm{x}\in\mathcal{X},\forall\mathcal{A}_{1},\mathcal{A}_{2}\in\mathcal{S}_{\mathcal{A}}$, where $\mathcal{X}$ is the input space.
+Based on this, we have the following theorem:
+Theorem 3.1.
+Given a perfectly-trained encoder $\mathcal{F}_{\theta}$ based on the augmentations sampled from predefined augmentation space $\mathcal{S}_{\mathcal{A}}$, it is impossible to inject a backdoor with trigger function $\mathcal{I}\in\mathcal{S}_{\mathcal{A}}$.
+The proof of this theorem can be found in the section A4.
+In practice, the boundary of the learned augmentation space for trained encoders $\mathcal{S}_{\mathcal{A}}^{\prime}$ is often imprecise, and it potentially reflects a relaxation of the predefined augmentation space.
+Consequently, using trigger functions that are not precisely within the predefined augmentation space $\mathcal{S}_{\mathcal{A}}$ but are instead distributionally close to it can also make achieving high attack success rates hard.
+Empirical Evidence.
+We also conduct experiments to confirm the invisible triggers designed for supervised learning are actually entangled with the inherent augmentations in self-supervised learning.
+Specifically, we use a ResNet18 pre-trained with SimCLR for a binary classification task to differentiate between samples poisoned by WaNet and those augmented by SimCLR.
+We ensure consistent feature representations for clean samples between the backdoored and clean models using utility loss from Jia et al [29].
+Results indicate that the models struggle to differentiate between the two categories.
+A t-SNE visualization of their features, as presented in Figure 3, indicates a significant overlap and entanglement.
+From this, we infer that the diminished effectiveness of supervised backdoor attack methods in SSL scenarios is attributed to the distributional similarity between the features of contrastive-learning-augmented samples and backdoor samples.
+The reason for the entanglement phenomenon on the invisible backdoor trigger is that the learned augmentation space reflects a relaxation of the predefined augmentation space, and such relaxation covers most of the invisible minor transformations.
+Thus, we aim to search for invisible transformations that can escape the inherent augmentation space.
 4 Attack Design
 4.1 Threat Model
+We follow the well-defined training-based backdoor injection threat model introduced in BadEncoder [27].
 Objectives of the Attacker.
-The objective of an attacker is to implant backdoors into a pre-trained image encoder by SSL (self-supervised learning).
+The objective of an attacker is to implant backdoors into a pre-trained image encoder by SSL.
 We define a backdoored image encoder model as $\mathcal{F}_{\theta}$ and the backdoor injector as $\mathcal{I}_{\phi}$.
@@
 $\mathcal{C}_{\epsilon}(\mathcal{F}_{\theta}(\boldsymbol{x}))=y,\quad\mathcal{C}_{\epsilon}(\mathcal{F}_{\theta}(\mathcal{I}_{\phi}(\boldsymbol{x})))=c$ (2)
-The ultimate aim of the attacker is to modify a clean image encoder and create a backdoored version that accomplishes several key objectives: ① Effectiveness: The backdoored model should achieve a high rate of successful attacks while preserving high accuracy in benign conditions.
-② Naturalness and Stealthiness: To avoid detection during human inspection, the sample with the trigger should appear authentic and natural.
-③ Robustness: The effectiveness of the attack should remain intact even when the triggered samples undergo common pre-processing steps, such as samples under various noises.
-Furthermore, the attack should bypass prevailing defense mechanisms.
+The attacker’s goal is to modify a clean image encoder to create a backdoored version that meets two key objectives: ① Effectiveness: The backdoored model should maintain a high attack success rate while preserving accuracy in benign conditions, keeping backdoored accuracy close to clean accuracy for downstream classifiers.
+② Naturalness and Stealthiness: The triggered sample should appear authentic and natural to avoid detection by human inspection.
 Attacker’s Knowledge and Capabilities.
-We assume that the attacker has access to a clean image encoder pre-trained following [29].
-Additionally, it is presumed that the attacker has access to a collection of unlabeled images, referred to as shadow dataset.
+Following BadEncoder [27], we assume that the attacker has access to a pre-trained clean image encoder and the attacker has full knowledge about the pre-trained encoder, such as the SSL method and the detailed contrastive augmentation operation used in pre-training.
+Additionally, it is presumed that the attacker can access a collection of unlabeled images, referred to as shadow dataset.
 The attacker is also assumed to have access to a few images from the Internet, called reference inputs, for each combination of a target downstream task and a target class.
-We assume that the attacker is capable of manipulating the training procedure to create an encoder with embedded backdoors.
-Accordingly, the attacker also has access to the augmentation transforms used to pre-train the encoder, which can be utilized in IMPERATIVE to generate stealthy and effective backdoor triggers.
+We assume that the attacker can manipulate the training procedure to create an encoder with embedded backdoors.
+Accordingly, the attacker also has access to the augmentation transforms used to pre-train the encoder, which can be utilized in INACTIVE to generate stealthy and effective backdoor triggers.
 However, we assume that the attacker cannot interfere with the training process of these downstream classifiers, such as the training dataset, model framework, and weights.
-4.2 Preliminaries
-HSV and HSL Color Spaces.
-HSV and HSL color spaces align more closely with human color perception, unlike RGB’s additive mixing.
-They allow for intuitive adjustments of tint, shade, and tone through single-dimension modifications.
-They have wide applications in the computer vision domain [62, 76, 42, 13].
-More details are shown in subsection D.5.
-Wasserstein Distance.
-Wasserstein distance is commonly employed for quantifying differences between two latent distributions without known common support or density functions [71].
-Specifically, we focus on the 2-Wasserstein distance, which is elaborated in subsection D.5.
+Unlike data-poisoning-based methods such as CTRL [34] and BLTO [63], our approach does not rely on matching distributions between pre-training and downstream datasets, allowing for broader applicability without interfering in downstream training data, models, or weights.
 4.2 Overarching Idea
-According to the observations in section 3, the key to enhancing the attack success rate in SSL is to disentangle the two overlapping distributions of the backdoor and augmentation transformation in the contrastive learning of the SSL’s pre-training stage.
-To widen the distributional gap, we initially convert the image from RGB color space to HSV and HSL color spaces, because they align better with the characteristics of augmentative transformations, such as “ColorJitter” (a data augmentation technique that randomly alters the brightness, contrast, saturation, and hue of an image).
+According to the previous observations, the key to enhancing the ASR in SSL is to disentangle the two overlapping distributions of the backdoor and augmentation transformation in the contrastive learning of the SSL’s pre-training stage.
+Gray Scaling and Color Jittering are necessary augmentations used in the self-supervised learning and most of existing SSL methods (e.g., SimCLR [6], MoCo [20], SimSiam [7] and BYOL [16]) use them.
+A detailed summary of the augmentation operations in different mainstream SSL methods can be found in Table A5.
+Since these augmentations primarily alter the color semantics of inputs, HSV and HSL color spaces serve as ideal input spaces for capturing and enlarging these effects [28].
+We aim to identify a trigger that escapes the inherent augmentation space of self-supervised learning by increasing the distance between backdoored samples and non-backdoored samples within the HSV and HSL color spaces.
 Since we already know the augmentation ways in the pre-training stage, thus we design $\mathcal{L}_{\text{disentangle}}$ to quantify the distributional gap between images in the two batches, which involves measuring the difference in color characteristics.
@@
 Hence we design $\mathcal{L}_{\text{stealthy}}$ to blend the backdoor seamlessly with the original image.
-4.4 Our Approach: IMPERATIVE
+4.3 Our Approach: INACTIVE
 In our context, we refer to a clean pre-trained image encoder and its backdoor-injected one as $\mathcal{F}_{\theta}$ and $\mathcal{F}_{\theta}^{\prime}$.
@@
 We denote a clean input $\boldsymbol{x}$ embedded with a trigger as $\boldsymbol{x}^{\prime}$, which is called a backdoored input.
-4.4.1 Enhancing Distributional Gap Between Backdoor Images and Augmented Images
+Enhancing Distributional Gap Between Backdoor Images and Augmented Images.
 To effectively enlarge the distributional distinction between the backdoor and the augmentation in SSL, we design $\mathcal{L}_{\text{disentangle}}$ for scenarios where augmentation transformations might weaken or obscure the pattern of backdoor triggers, leading to a decrease in the ASR.
@@
 The disentangle loss is defined as follows.
-$\begin{array}[]{rl}\mathcal{L}_{\text{disentangle}}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum\limits_{\boldsymbol{x}\in D_{s}}&\|H\left(\boldsymbol{x}^{\prime}\right)-H(\tilde{\boldsymbol{x}})\|_{2}+\|S\left(\boldsymbol{x}^{\prime}\right)-S(\tilde{\boldsymbol{x}})\|_{2}\\ +&\|V\left(\boldsymbol{x}^{\prime}\right)-V(\tilde{\boldsymbol{x}})\|_{2}+\|L\left(\boldsymbol{x}^{\prime}\right)-L(\tilde{\boldsymbol{x}})\|_{2},\end{array}$ (2)
-where $H$,$S$,$V$,$L$ denote Hue, Saturation, Value, and Lightness transformations from HSV and HSL color spaces, which are elaborated on in subsection 4.2.
+$\begin{array}[]{rl}\mathcal{L}_{\text{disentangle}}=\\ -\frac{1}{\left|\mathcal{D}_{s}\right|}\sum\limits_{\boldsymbol{x}\in D_{s}}&\|H\left(\boldsymbol{x}^{\prime}\right)-H(\tilde{\boldsymbol{x}})\|_{2}+\|S\left(\boldsymbol{x}^{\prime}\right)-S(\tilde{\boldsymbol{x}})\|_{2}\\ +&\|V\left(\boldsymbol{x}^{\prime}\right)-V(\tilde{\boldsymbol{x}})\|_{2}+\|L\left(\boldsymbol{x}^{\prime}\right)-L(\tilde{\boldsymbol{x}})\|_{2},\end{array}$ (3)
+where $H$,$S$,$V$,$L$ denote Hue, Saturation, Value, and Lightness from HSV and HSL color spaces.
 We denote an input $\boldsymbol{x}$ augmented by the transformations used in the encoder’s pre-training stage as $\tilde{\boldsymbol{x}}$.
@@
 $\|u-v\|_{2}$ denotes the $\ell_{2}$ distance between sample $u$ and sample $v$.
-4.4.2 Feature Alignment Between Backdoored and Reference Images
-Follwoing BadEncoder [30], we enhance the backdoor attack effectiveness by making the compromised image encoder outputs similar feature embeddings for any sample injected with backdoor $\boldsymbol{x}^{\prime}$ in the shadow dataset $\mathcal{D}_{s}$ and the reference inputs $\mathcal{R}_{i}$ of a pair $\left(T_{i},y_{i}\right)$.
+Feature Alignment Between Backdoored and Reference Images.
+Following BadEncoder, we enhance the backdoor attack effectiveness by making the compromised image encoder outputs similar feature embeddings for any sample injected with backdoor $\boldsymbol{x}^{\prime}$ in the shadow dataset $\mathcal{D}_{s}$ and the reference inputs $\mathcal{R}_{i}$ of a pair $\left(T_{i},y_{i}\right)$.
 Consequently, a compromised downstream classifier developed from our compromised image encoder is inclined to assign identical labels to both reference samples $\mathcal{R}_{i}$ and to any compromised sample $\boldsymbol{x}^{\prime}$.
-We call this process feature alignment between backdoored and reference images, and the $\mathcal{L}_{alignment}$ is defined as follows.
-$\mathcal{L}_{alignment}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}^{\prime}\right),\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right)\right)}{\left|\mathcal{D}_{s}\right|\cdot\sum_{i=1}^{t}r_{i}},$ (3)
+We call this process feature alignment between backdoored and reference images, and the $\mathcal{L}_{\text{alignment}}$ is defined as follows.
+$\mathcal{L}_{\text{alignment}}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}^{\prime}\right),\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right)\right)}{\left|\mathcal{D}_{s}\right|\cdot\sum_{i=1}^{t}r_{i}},$ (4)
 where $s(\cdot,\cdot)$ is used to quantify the degree of similarity, for instance, cosine similarity, between a pair of feature embeddings.
 The term $|D_{s}|$ denotes the count of samples within the shadow dataset, and the denominators serve the purpose of standardizing the losses.
-4.4.3 Preserving Covert and Natural Backdoors with Expanded Distributional Gaps
+Preserving Covert and Natural Backdoors with Expanded Distributional Gaps.
 We employ several metrics that measure the similarity between the backdoored image and the original one in both pixel and feature space to ensure that our trigger remains both natural and inconspicuous.
-For assessing similarity in pixel space, we use SSIM [89] and PSNR [26].
-Meanwhile, for high-level feature space comparisons, we first use LPIPS [96] which better reflects the subjective experience of image quality and similarity, capturing nuances that simpler metrics might miss.
-Moreover, since our objective is to reduce the distributional disparity between backdoored and clean samples, the 2-Wasserstein distance (WD) [82] serves as an effective tool for this purpose.
-We first extract features from backdoored and clean samples respectively using a pre-trained model $\mathcal{M}$, such as ResNet34 [21].
-Then we use WD to measure the distance between the extracted features.
-$\mathcal{L}_{\text{stealthy }}$ is defined as follows.
-$\begin{array}[]{rl}\mathcal{L}_{\text{stealthy }}=\sum_{\boldsymbol{x}\in D_{s}}&\lambda_{1}\cdot\mathrm{LPIPS}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right)+\mathrm{WD}\left(\mathcal{M}(\boldsymbol{x}^{\prime}),\mathcal{M}(\boldsymbol{x})\right)\\ -&\lambda_{2}\cdot\mathrm{PSNR}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right)-\mathrm{SSIM}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right),\end{array}$ (4)
+To assess similarity in pixel space, we use SSIM and PSNR.
+Meanwhile, for high-level feature space comparisons, we first use LPIPS which better reflects the subjective experience of image quality and similarity.
+Following [82] and [68], we also use Wasserstein distance [71] (WD) to reduce the distributional disparity between backdoored and clean samples.
+$\mathcal{L}_{\text{stealthy }}$ is defined as follows:
+$\begin{array}[]{rl}\mathcal{L}_{\text{stealthy }}=\\ \sum_{\boldsymbol{x}\in D_{s}}&\lambda_{1}\cdot\mathrm{LPIPS}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right)+\mathrm{WD}\left(\mathcal{M}(\boldsymbol{x}^{\prime}),\mathcal{M}(\boldsymbol{x})\right)\\ -&\lambda_{2}\cdot\mathrm{PSNR}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right)-\mathrm{SSIM}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right),\end{array}$ (5)
 where $\lambda_{1},\lambda_{2}$ are used to scale different loss terms to the same scale from 0 to 1.
-Algorithm 1 Our backdoor attack IMPERATIVE
-Input: Pre-trained benign encoder $\mathcal{F}_{\theta}^{*}$, shadow dataset $\mathcal{D}_{s}$, reference input set $\mathcal{R}$
+Algorithm 1 Our backdoor attack INACTIVE
+Input: Pre-trained clean encoder $\mathcal{F}_{\theta}^{*}$, shadow dataset $\mathcal{D}_{s}$, reference input set $\mathcal{R}$
 Output: Backdoored encoder $\mathcal{F}_{\theta}$ , backdoor trigger injector $\mathcal{I}_{\phi}$
@@
 2:    $\mathcal{F}_{\theta}\leftarrow\mathcal{F}_{\theta}^{*}$; $\hat{\mathcal{D}_{s}}\leftarrow$ Augment samples in $\mathcal{D}_{s}$;
-3:     ϕsubscriptitalic-$\mathcal{I}_{\phi}\leftarrow$ a pre-trained backdoor injector $\mathcal{I}_{\phi}^{*}$ using Algorithm 2
+3:    ϕsubscriptitalic-$\mathcal{I}_{\phi}\leftarrow$ a pre-trained backdoor injector $\mathcal{I}_{\phi}^{*}$ using Alg. 2
 4:    for $\text{iter}=0$ to max_epochs do
@@
 8:       $\triangleright$ Equation 5
-9:         $\mathcal{L}_{alignment}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}\left(\boldsymbol{x}^{\prime}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\left|\mathcal{D}_{s}\right|\cdot\sum_{i=1}^{t}r_{i}}$triangleright Equation 3
+9:       $\mathcal{L}_{\text{alignment}}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}\left(\boldsymbol{x}^{\prime}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\left|\mathcal{D}_{s}\right|\cdot\sum_{i=1}^{t}r_{i}}$triangleright Equation 4
 10:       $\mathcal{L}_{\text{injector}}=\mathcal{L}_{\text{stealthy}}+\alpha\cdot\mathcal{L}_{\text{disentangle}}+\beta\cdot\mathcal{L}_{\text{alignment}}$triangleright Equation 6
 11:       $\phi_{\mathcal{I}}=\phi_{\mathcal{I}}-lr_{1}\cdot\frac{\partial\mathcal{L}_{\text{injector}}}{\partial\phi_{\mathcal{I}}}$
-12:         $\mathcal{L}_{\text{consistency }}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\sum_{i=1}^{t}r_{i}}$triangleright Equation 6
-13:         $\mathcal{L}_{\text{utility}}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}(\boldsymbol{x}),\mathcal{F}_{\theta}(\boldsymbol{x})\right)$$\triangleright$ Equation 7
+12:       $\mathcal{L}_{\text{consistency }}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\sum_{i=1}^{t}r_{i}}$triangleright Equation A1
+13:       $\mathcal{L}_{\text{utility}}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}(\boldsymbol{x}),\mathcal{F}_{\theta}(\boldsymbol{x})\right)$$\triangleright$ Equation A2
 14:       $\mathcal{L}_{\text{encoder}}=\mathcal{L}_{\text{alignment}}+\mathcal{L}_{\text{consistency}}+\mathcal{L}_{\text{utility}}$
 15:       $\theta_{\mathcal{F}}=\theta_{\mathcal{F}}-lr_{2}\cdot\frac{\partial\mathcal{L}_{\text{encoder}}}{\partial\theta_{\mathcal{F}}}$
-4.4.4 Optimization Problem Formulation and Algorithm
-We have already defined $\mathcal{L}_{\text{disentangle}}$, $\mathcal{L}_{\text{stealthy }}$, $\mathcal{L}_{alignment}$ in the sections above.
-Then we can define our IMPERATIVE as an optimization problem.
+Algorithm 2 Pre-training backdoor injector
+Input: Shadow dataset $\mathcal{D}_{s}$       Output: Pre-trained backdoor injector $\mathcal{I}_{\phi}$
+1:function Pre-training injector($\mathcal{D}_{s}$)
+2:    $\hat{\mathcal{D}_{s}}\leftarrow$ Augment samples in $\mathcal{D}_{s}$; $\mathcal{I}_{\phi}\leftarrow\text{Random initialization}$
+3:    for $\text{iter}=0$ to max_epochs do
+4:       ${\mathcal{D}_{s}}^{{}^{\prime}}\leftarrow\mathcal{I}_{\phi}(\mathcal{D}_{s})$
+5:       $\mathcal{L}_{\text{disentangle}}\leftarrow$ distribution difference between backdoor images $\boldsymbol{x}^{\prime}$ and augmented images $\tilde{\boldsymbol{x}}$, $\forall\boldsymbol{x}^{\prime}\in{\mathcal{D}_{s}}^{{}^{\prime}},\forall\tilde{\boldsymbol{x}}\in\hat{\mathcal{D}_{s}}$triangleright Equation 3
+6:       $\mathcal{L}_{\text{stealthy}}\leftarrow$ distance between backdoor image $\boldsymbol{x}^{\prime}$ and clean image $\boldsymbol{x}$, $\forall\boldsymbol{x}^{\prime}\in{\mathcal{D}_{s}}^{{}^{\prime}},\forall\boldsymbol{x}\in\mathcal{D}_{s}$triangleright Equation 5
+7:       $\mathcal{L}_{\text{ours}}=\mathcal{L}_{\text{stealthy}}+\mu\cdot\mathcal{L}_{\text{disentangle}}$
+8:       $\phi_{\mathcal{I}}=\phi_{\mathcal{I}}-lr\cdot\frac{\partial\mathcal{L}_{\text{ours}}}{\partial\phi_{\mathcal{I}}}$
+Optimization Problem Formulation and Algorithm.
+We have defined $\mathcal{L}_{\text{disentangle}}$, $\mathcal{L}_{\text{stealthy }}$, $\mathcal{L}_{\text{alignment}}$ in the sections above.
+Then we can define our INACTIVE as an optimization problem.
 Concretely, our backdoor trigger injector $\mathcal{I}_{\phi}$ is a solution to the subsequent optimization problem:
-$\min_{\mathcal{I}_{\phi}}\mathcal{L}=\mathcal{L}_{\text{stealthy }}+\alpha\cdot\mathcal{L}_{\text{disentangle}}+\beta\cdot\mathcal{L}_{\text{alignment}},$ (5)
+$\min_{\theta_{\mathcal{F}}}\min_{\mathcal{I}_{\phi}}\mathcal{L}=\mathcal{L}_{\text{stealthy }}+\alpha\cdot\mathcal{L}_{\text{disentangle}}+\beta\cdot\mathcal{L}_{\text{alignment}},$ (6)
 where $\alpha$ and $\beta$ serve as hyper-parameters to provide equilibrium among these three loss components.
-We adopt Algorithm 1 to solve the optimization problem, where we alternatively optimize the backdoor injector and the compromised image encoder, and output the final backdoored encoder $\mathcal{F}_{\theta}$ and backdoor trigger injector $\mathcal{I}_{\phi}$.
-Additionally, to speed up the optimization process and promote the backdoor attack efficacy, we adopt Algorithm 2 to pre-train a backdoor injector to initialize the injector in Algorithm 1 (Please refer to Appendix D.2).
-We further conduct ablation study to prove the function of the injector pre-training in Appendix B.
-Finally, we prove that the performance of our model is not sensitive to the hyperparameters in the Appendix C.2.
-4.4.5 Highlight of Our Technical Contributions
-We wish to highlight our distinct technical contributions: ①Identifying the reason behind existing triggers’ ineffectiveness: We discover that existing backdoor attacks on SSL, such as BadEncoder [30], attain high ASRs but utilize visible triggers.
-Moreover, traditional imperceptible triggers designed for supervised classifiers, e.g., WaNet [51], ISSBA [41], do not perform effectively in SSL, owing to the overlap in feature-space distributions between the backdoor samples and the augmented samples used in SSL models.
-② Proposing disentanglement strategy to generate stealthy and effective backdoor triggers: Based on our insights, we propose a novel backdoor attack in SSL that effectively maintains the potency, stealthiness, and robustness of the generated trigger by disentangling the distribution of backdoor samples from augmented samples within SSL.
+We adopt Alg.
+1 to solve the optimization problem, where we alternatively optimize the backdoor injector and the compromised image encoder and output the final backdoored encoder $\mathcal{F}_{\theta}$ and backdoor trigger injector $\mathcal{I}_{\phi}$.
+Additionally, to speed up the optimization process and promote the backdoor attack efficacy, we adopt Alg.
+2 to pre-train a backdoor injector to initialize the injector in Alg.
+1.
+We use the U-Net architecture [54] for the backdoor injector, as shown in the Table A12.
 5 Evaluation
-In this section, we first evaluate the effectiveness and stealthiness of IMPERATIVE using four datasets in subsection 5.2 and subsection 5.3.
-Then we evaluate the robustness of our method against various SOTA backdoor defenses and noises of different depths in subsection 5.4 and Appendix C.1.
-After that, we prove our method’s efficacy in large-scale dataset in subsection 5.5.
-Furthermore, to prove its generalization capability, we conduct more attacking cases on various representative SSL algorithms and a multi-modal model with various augmentations in Appendix A.
-Moreover, we conduct the ablation study in Appendix B to prove each component’s function.
-Finally, we evaluate the sensitivity of parameters in Appendix C.2 and further analyze the performance of our method in Appendix C.3.
+We first evaluate the effectiveness and stealthiness of INACTIVE using four datasets, followed by an assessment of its robustness against various backdoor defenses and noises.
+To demonstrate generalization, we conduct additional attacks on various SSL algorithms and a multi-modal model with different augmentations, detailed in section A1 and section A2.
+section A3 further validates each component’s role.
+section A5 examines parameter sensitivity and performance.
+Pre-training Dataset Downstream Dataset No Attack BadEncoder + WaNet trigger BadEncoder + CTRL trigger BadEncoder + Ins-Kelvin trigger BadEncoder + Ins-Xpro2 trigger DRUPE + Patch trigger Ours
+[…]
+STL10 CIFAR10 86.77 84.43 10.28 87.19 8.72 86.75 18.63 86.85 16.83 84.36 98.39 87.11 99.58
+GTSRB 76.12 74.45 5.23 77.57 8.17 76.49 72.95 76.71 14.02 75.93 96.09 75.82 97.97
+SVHN 55.35 58.29 16.83 54.29 3.32 56.67 38.03 58.42 18.68 75.64 96.68 58.62 99.76
+CIFAR10 STL10 76.14 72.73 9.78 75.73 16.85 74.89 1.16 74.11 5.91 74.43 96.72 74.02 99.68
+GTSRB 81.84 75.85 5.46 79.94 97.95 78.56 2.50 75.08 42.40 80.35 97.22 79.15 98.73
+SVHN 61.52 54.79 17.99 66.33 40.91 68.49 22.13 68.95 30.91 76.02 96.23 63.67 98.79
+Average / 72.96 70.09 10.93 73.51 29.32 73.64 25.90 73.35 21.46 77.79 96.89 73.10 99.09
+Table 1: Effectiveness comparison to representative backdoor attacks in SSL with different triggers (CA(%), BA(%), and ASR(%)).
+We compare our method to BadEncoder [27] with various existing stealthy triggers.
+We also include the results of DRUPE [68] with their default visible patch trigger.
+We include CTRL here to demonstrate that it is ineffective across various downstream datasets.
+Our approach constantly achieves the highest ASRs while maintaining the accuracy on clean samples of the downstream classifiers trained on the backdoored encoder.
 5.1 Experimental Setup
 Datasets.
-We utilize the following four image datasets, i.e.
-CIFAR10 [34], STL10 [9], GTSRB [68], SVHN [50] to evaluate our method, which are also frequently utilized in research related to backdoor attacks [30, 11, 16, 51].
-We elaborate on the details of the datasets in Appendix D.6.1.
-Pre-training Dataset Downstream Dataset WaNet CTRL Ins-kelvin Ins-xpro2 Ours
+We utilize four image datasets, i.e.
+CIFAR10 [31], STL10 [9], GTSRB [62], SVHN [46] and ImageNet [55] to evaluate our method, which are also frequently used in backdoor attacks research [27, 47].
+More details are introduced in subsection A6.5.
+Evaluation Metrics.
+To assess the effectiveness of our method, we employ three metrics following existing works [88, 27]: Clean Accuracy (CA): the accuracy of a clean downstream classifier on clean testing images from the downstream dataset; Benign Accuracy (BA): the accuracy of a backdoored downstream classifier on the same clean testing images from the downstream dataset; Attack Success Rate (ASR): the success rate of backdoor attacks.
+To evaluate the stealthiness and naturalness of the backdoor triggers, we employ three metrics following existing works [28]: SSIM [73], PSNR [25], LPIPS [87], Feature Similarity Indexing Method (FSIM) [86] and Fréchet Inception Distance(FID) [21].
+Higher SSIM, PSNR, FSIM and lower LIPIPS, FID indicate better stealthiness and naturalness of the generated backdoored images.
+SSL Frameworks.
+In the pre-training stage, we employ SimCLR [6] by default to train a ResNet18 [19] model, serving as our image encoder.
+Furthermore, we prove the effectiveness of our method on other SSL frameworks, i.e., MoCo [20], BYOL [16], SimSiam [7], SwAV [4], and CLIP [53] in section A1 and section A2.
+Attack Baselines.
+We select two Instagram filters, Kelvin and Xpro2, as baseline triggers for aesthetic enhancements [30, 42].
+Additionally, WaNet [47], CTRL [34], and ISSBA [36] are chosen for their stealthiness and high ASR.
+These triggers are injected into compromised encoders using BadEncoder.
+We also include DRUPE [68], a SOTA backdoor method using SimCLR and a patch trigger, as a baseline.
+To ensure a fair comparison, we evaluate our method against CTRL [34], SSLBKD [56], POIENC [39], and BLTO [63] using the same CIFAR10 as the pre-trained and downstream dataset under SimCLR, BYOL, and SimSiam.
+For SSLBKD, the trigger is randomly placed, while for SSLBKD-fixed, it’s in the lower-right corner.
+We show more experimental settings and details in subsection A6.5.
+Attack Invisible SSL Method
+SimCLR BYOL SimSiam
 […]
-STL10 CIFAR10 86.77 84.43 10.28 87.19 8.72 86.75 18.63 86.85 16.83 87.11 99.58
-GTSRB 76.12 74.45 5.23 77.57 8.17 76.49 72.95 76.71 14.02 75.82 97.97
-SVHN 55.35 58.29 16.83 54.29 3.32 56.67 38.03 58.42 18.68 58.62 99.76
-CIFAR10 STL10 76.14 72.73 9.78 75.73 16.85 74.89 1.16 74.11 5.91 74.48 95.00
-GTSRB 81.84 75.85 5.46 79.94 97.95 78.56 2.50 75.08 42.40 79.15 98.73
-SVHN 61.52 54.79 17.99 66.33 40.91 68.49 22.13 68.95 30.91 63.67 98.79
-Table 1: Results of CA(%), BA(%), and ASR(%) in comparison experiments. Our IMPERATIVE achieves high ASRs while maintaining the accuracy of the downstream classifiers.
-Evaluation Metrics.
-To assess the effectiveness of our IMPERATIVE, we employ three metrics following existing works [97, 30]: Clean Accuracy (CA), Backdoored Accuracy (BA), Attack Success Rate (ASR).
-More details are explained in Appendix D.6.2.
-To evaluate the stealthiness and naturalness of our IMPERATIVE, we employ three metrics following existing works [31, 84]: SSIM [89], PSNR [26] and LPIPS [96].
-In experiments, they measure the similarity between backdoored images and original clean images.
-Higher SSIM and PSNR and lower LIPIPS indicate better stealthiness and naturalness of the generated backdoored images.
-Parameter Setting.
-We consider a scenario where the attacker chooses a specific downstream task/dataset, a specific target class, and a specific reference input.
-For the datasets CIFAR10, STL10, GTSRB, and SVHN, the selected target classes are “airplane”, “truck”, “priority sign”, and “digit one”, respectively.
-CIFAR10 and STL10 are used as the default pre-training dataset.
-In the pre-training stage, we employ SimCLR [5] by default to train a ResNet18 [21] model, serving as our image encoder.
-Furthermore, we prove the effectiveness of our method across various representative SSL frameworks in Appendix A, including MoCo [22], BYOL [17], SimSiam [6], DINO [4], SwAV [3], and CLIP [58].
-More details are shown in Appendix D.6.3.
-Baselines.
-We first choose two normal Instagram filters, i.e., kelvin and xpro2, as our baseline triggers, which are gradient-free methods to apply aesthetic enhancements to images, such as adjusting color balance, contrast, and saturation [33, 46].
-Moreover, WaNet [51], and CTRL [37], ISSBA [41] are also chosen as our baseline triggers to compare with our method.
-They are state-of-the-art backdoor triggers and outperform earlier backdoor triggers in terms of stealthiness, yet maintain a high ASR.
-They are implemented using the default hyperparameters as specified in their original papers.
-Note that we use BadEncoder [30] to inject these baseline triggers into the compromised encoders.
+POIENC [39] 80.50 11.10 81.70 10.70 81.90 10.70
+SSLBKD [56] 79.40 33.20 80.30 46.20 80.60 53.10
+SSLBKD (fixed) [56] 80.00 10.50 82.30 11.20 81.90 10.70
+CTRL [34] 80.50 85.30 82.20 61.90 82.00 74.90
+BLTO [63] 90.10 91.27 91.21 94.78 90.18 84.63
+Ours 90.19 100.00 93.01 99.99 91.01 99.99
+Table 2: Effectiveness comparison to data-poisoning-based backdoor attacks in SSL with their default triggers.
+We show the results of BA(%), and ASR(%) with the same pre-trained and downstream dataset CIFAR10.
+Since data poisoning-based methods require matched distributions between pre-training and downstream, we use the same pre-trained and downstream datasets.
+Our threat model is different from theirs, and our method can be applied when the distributions of pre-training and downstream datasets are different.
+This table’s key aim is to demonstrate that our method achieves much higher ASR than them.
 5.2 Effectiveness Evaluation
-Effective Attack.Table 1 indicates our method consistently achieves high ASRs (most above 98%) across various datasets.
-For instance, with STL10 as the pre-training dataset and CIFAR10 as the downstream dataset, our method reaches an ASR of 99.58%, indicating a highly successful rate of triggering the desired misclassification.
-This is significantly higher than the ASRs of other benchmark methods like WaNet [51] and CTRL [37] in the same context.
-Accuracy Preservation.Table 1 proves that IMPERATIVE preserves the accuracy of the downstream classifiers.
-The differences between BA and CA are within 2% across the different combinations of pre-training and downstream datasets.
-This suggests that the presence of the backdoor introduced by our method does not compromise the classifier’s ability to correctly label clean images.
-For instance, when STL10 is used for pre-training and GTSRB is the downstream dataset, the BA is 76.12% compared to the CA of 75.82%, showing a negligible impact on the classifier’s performance.
+Effective Attack.
+As shown in Table 1, with different pre-trained and downstream datasets, our method achieves a high average ASR of 99.09% across various datasets.
+Additionally, Table 2 demonstrates that with the same pre-trained and downstream datasets, our approach also achieves nearly 100% ASRs.
+Our method outperforms all baseline methods in all scenarios, highlighting its robustness and superior effectiveness in executing successful backdoor attacks.
+Accuracy Preservation.
+The downstream classifiers trained on the backdoored encoder maintain good accuracy on clean samples, as shown in Table 1.
+The average BA is 73.10% compared to the average CA of 72.96%, with the difference within 1%.
+This suggests that the backdoor introduced by our method does not compromise the classifier’s ability to label clean images correctly.
+This is because $\mathcal{L}_{\text{utility}}$ guarantees that the backdoored and clean image encoders yield similar feature vectors for clean inputs.
+5.3 Stealthiness Evaluation
 […]
-5.3 Stealthiness Evaluation
-As shown in Table 3, our method exhibits strong stealthiness advantages with SSIM values consistently close to 1 and most above 0.95, indicating minimal structural changes to the images.
-PSNR values are notably high, exceeding 40 for some cases, which implies that the noise introduced is hardly detectable.
-Furthermore, the LPIPS values are exceptionally low, ranging from 0.0001 to 0.02, demonstrating that the perceptual difference between the original and perturbed images is almost negligible.
-These figures collectively suggest that our method introduces changes that are virtually imperceptible to both human eyes and standard image quality metrics.
+Badencoder [27]/DRUPE [68] 0.8355 14.1110 0.07693 0.820 53.363
+CTRL [34] 0.9025 32.4098 0.00034 0.865 71.138
+WaNet [47] 0.7704 14.2372 0.07432 0.662 98.092
+Ins-Kelvin [42] 0.4955 16.1925 0.14000 0.677 96.449
+Ins-Xpro2 [42] 0.5981 17.9173 0.04434 0.817 35.084
+POIENC [39] 0.1214 11.2787 0.15867 0.597 172.220
+SSLBKD [56] 0.8737 16.2414 0.09640 0.891 118.320
+BLTO [63] 0.8417 29.6756 0.00941 0.950 36.385
+Ours 0.9633 35.8649 0.00896 0.969 16.320
+Table 3: Stealthiness comparison to existing methods on CIFAR10. Our method remains stealthy. Detailed data are shown in Table A9.
+Algorithmic Metrics.
+We first compare the average SSIM, PSNR, and LPIPS when the pre-trained dataset is CIFAR10 and downstream datasets are STL10, GTSRB, and SVHN injected with these backdoor triggers to compare the stealthiness of various backdoor attack methods.
+Table 3 indicates that our method exhibits strong stealthiness advantages with an average of 0.9633 SSIM, 35.8649 PSNR, 0.00896 LIPIS, 0.969 FSIM, and 16.320 FID indicating minimal structural changes to the images, hardly detectable noise, almost negligible perceptual difference between the original and perturbed images.
+Although CTRL achieves a better LPIPS, our method outperforms it in both SSIM and PSNR.
+Additionally, our average ASR is 99.09%, significantly higher than CTRL 29.32%(see Table 1), indicating that our method is more effective overall.
+More detailed data across various datasets, i.e., CIFAR10, STL10, GTSRB, SVHN, and ImageNet are shown in Table A9 and Table A10.
 5.4 Robustness Evaluation
-To assess the resilience of our method against current backdoor defenses, we deploy various state-of-the-art backdoor defense strategies, i.e., DECREE [15], Beatrix [48], Neural Cleanse [83], STRIP [16], Grad-CAM [64] and measure how effectively our method withstands these defenses.
-Additionally, to further test the robustness of our method, we evaluate its endurance against the following commonly studied noises, i.e., JPEG compression [10, 12, 20, 14], Poisson noise [1, 91], and Salt&Pepper noise [38, 1, 91].
-We show the defense results of STRIP, Grad-CAM, and noises in Appendix C.1.
+To assess the resilience of our method against current backdoor defenses, we deploy various SOTA backdoor defense strategies, i.e., DECREE [13], Beatrix [44], ASSET [48], Neural Cleanse (NC) [72], STRIP [14], Grad-CAM [58] for evaluation.
+Additionally, to further test the robustness of our method, we evaluate its endurance against the following commonly studied noises, i.e., JPEG compression [10, 11], Poisson noise [1, 79], and Salt&Pepper noise [35, 1].
+We also design an adaptive defense method for INACTIVE.
+We show that INACTIVE cannot be defended by STRIP, NC, Grad-CAM, noises, and adaptive defense in subsection A5.1.
 DECREE.
-[15]  We initially assess whether our method can elude detection by DECREE [15], a method for identifying Trojan attacks in pre-trained encoders.
-It judges an encoder to be compromised if the proportion of the reversed trigger’s $\mathcal{L}^{1}$ norm relative to the greatest $\mathcal{L}^{1}$ norm within the input space of the model falls below a threshold, i.e., 0.1 $\mathcal{P}\mathcal{L}^{1}$ Norm [15].
-As shown in Table 3, $\mathcal{P}\mathcal{L}^{1}$-Norm of each pre-trained and downstream dataset pair are above a default threshold 0.1, and thus this approach does not succeed in identifying backdoored encoders created by IMPERATIVE.
-This is because the stealthy loss is designed to narrow the distribution gap between backdoored and normal data in both pixel and feature levels.
-This makes it challenging for DECREE to detect abnormal model internal representation changes when specific behaviors are activated.
+DECREE [13] identifies trojan attacks in pre-trained encoders by flagging an encoder as compromised if the reversed trigger’s $\mathcal{L}^{1}$ norm proportion falls below a 0.1 threshold.
+As shown in Table 4, the $\mathcal{P}\mathcal{L}^{1}$-Norm for each pre-trained and downstream dataset pair exceeds this threshold, so DECREE fails to detect backdoored encoders created by INACTIVE.
+This is because our invisible trigger breaks DECREE’s assumption of a visible patch trigger, and our stealthy loss further narrows the distribution gap between backdoored and normal data, masking internal model anomalies.
+Pre-trained Dataset Downstream Dataset $\mathcal{P}\mathcal{L}^{1}$-Norm
+CIFAR10 STL10 0.25
+SVHN 0.39
+GTSRB 0.15
+STL10 CIFAR10 0.21
+SVHN 0.34
+GTSRB 0.20
+Table 4: Evaluation results of DECREE [13]. A model is judged as backdoored if its $\mathcal{P}\mathcal{L}^{1}$-Norm <0.1.
 Beatrix.
-[48]  Beatrix [48] identifies poisoned samples by detecting abnormalities in the feature space.
-We used two pretraining datasets, CIFAR-10 and STL-10, and created backdoored encoders using BadEncoder and IMPERATIVE.
+Beatrix [44] identifies poisoned samples by detecting abnormalities in the feature space.
+We use two pretraining datasets, CIFAR-10 and STL-10, and create backdoored encoders using BadEncoder and INACTIVE.
 By sampling 500 clean inputs and 500 poisoned samples, we applied Beatrix to differentiate them.
-The findings, as illustrated in Table 5, show that Beatrix effectively recognizes poisoned samples from BadEncoder with over 93% accuracy.
-However, Beatrix struggles to identify poisoned samples from IMPERATIVE, with a detection accuracy of below 50% on both CIFAR-10 and STL-10, which is like random guessing.
-We further analyze the reason why Beatrix fails to detect IMPERATIVE in Appendix C.1.1.
+We find (see Table 5) that Beatrix effectively recognizes poisoned samples from BadEncoder with over 93% accuracy.
+However, Beatrix struggles to identify poisoned samples from INACTIVE, with a detection accuracy of below 50% on both CIFAR-10 and STL-10, which is like random guessing.
+We further analyze the reasons for the defense failure in subsection A5.1.
+Encoder Method TP FP FN TN Acc
+CIFAR-10 BadEncoder 499 24 1 476 97.50%
+Ours 0 24 500 476 47.60%
+STL-10 BadEncoder 458 24 42 476 93.40%
+Ours 5 24 495 476 48.10%
+Table 5: Detection results by Beatrix [44]. It struggles to detect poisoned samples from ours.
+ASSET.
+ASSET aims to distinguish between backdoored and clean samples by eliciting distinct behaviors in the model when processing these two data types, facilitating their separation [48].
+We replicate their defensive techniques on our backdoored CIFAR-10 dataset.
+Specifically, we applied our synthesized trigger to CIFAR-10 (with a target label of 0) to create a poisoned version of CIFAR-10, maintaining a 100% poisoning rate as our default setting.
+The feature extractor used is the ResNet18 backbone, trained on this poisoned CIFAR-10 dataset.
+The True Positive Rate (TPR) measures how effectively a backdoor detection method identifies backdoored samples, with a higher TPR (closer to 100%) indicating stronger filtering capability.
+The False Positive Rate (FPR) reflects the precision of this filtering: when TPR is sufficiently high, FPR shows the trade-off, highlighting the proportion of clean samples incorrectly flagged as backdoored.
+A lower FPR suggests fewer clean samples are mistakenly discarded, ensuring more clean data is retained for further use.
+Based on ASSET’s metrics, we calculated the TPR as 7.14% and the FPR as 1.8%, indicating that our poisoned data can largely evade ASSET’s detection.
+5.5 Generalization to Large-scale Dataset
+We assess the generalization of our method on a large-scale dataset by attacking an ImageNet-pre-trained encoder from Google [6].
+We compare our method’s performance with ISSBA, which is also trained and tested on ImageNet in its paper.
+Experimental setups are detailed in subsection A6.5.
+Experimental Results.
+Table 6 indicates that our method is highly effective on ImageNet, with an average 98.66% ASR across different datasets.
+Moreover, Table 7 indicates the high SSIM and PSNR values and low LPIPS values, demonstrating that the perturbations made by INACTIVE are almost imperceptible.
+Moreover, the average 83.91% BA is close to the average 83.59% CA, indicating our attack maintains accuracy for the given downstream task despite the backdoor.
+Additionally, both our ASR and BA are much higher than those of the baseline ISSBA [36], proving ours has better performance.
+Downstream Dataset No Attack ISSBA [36] Ours
 […]
-Neural Cleanse.
-[83]  We then test the resilience of our method against a well-known reverse engineering defense, namely Neural Cleanse (NC) [83].
-Specifically, it utilizes the anomaly index to measure the deviation of the reconstructed triggers by their sizes, labeling models with an anomaly index exceeding two as Trojan-infected.
-Since NC is specifically designed for classifiers and cannot be directly used on image encoders, NC is employed to identify backdoors in a compromised downstream classifier.
-Moreover, the testing dataset from a targeted downstream dataset serves as the clean dataset.
-Experimental outcomes, as illustrated in Table 3, indicate that NC is unable to detect the Trojan model created by our approach.
-While effective in identifying patch-based Trojans [44], this method presumes a uniform trigger pattern at the pixel level across different samples.
-Our method circumvents this by employing input-dependent triggers, meaning the pixel-level Trojan alterations vary across samples.
-5.5 Generalization to Large-scale Dataset
-We the generalization ability of our method to large-scale dataset by applying our attack to an image encoder pre-trained on ImageNet [60] released by Google [5].
-We show the detailed experimental setup in Appendix D.6.4.
-Table 5 indicates that our method is highly effective on ImageNet.
-It achieves high ASR across different datasets, ranging from 96% to 100%, showing its capability to consistently deceive the target system.
-As shown in Table 3, the high SSIM values ranging from 0.989 to 0.998, high PSNR values ranging from 32.14 to 38.34, and low LPIPS values demonstrate that the perturbations made by IMPERATIVE are almost imperceptible, maintaining the visual quality of the images.
-Moreover, the BA is close to CA, indicating our attack maintains accuracy for the given downstream task despite the backdoor.
-Additionally, both our ASR and BA are much higher than those of the benchmark baseline ISSBA [41], proving ours has better performance.
+STL10 95.68 92.58 9.97 93.48 100.00
+GTSRB 80.32 66.29 5.10 82.84 96.00
+SVHN 74.77 67.67 18.03 75.40 99.99
+Average 83.59 75.51 11.03 83.91 98.66
+Table 6: Comparative results (CA(%), BA(%), and ASR(%)) of ISSBA [36] and our attack on ImageNet.
+Ours constantly achieves the highest ASRs while maintaining accuracy on clean samples of the downstream classifiers.
+[…]
+ISSBA [36] 0.7329 31.3496 0.12424
+Ours 0.9867 34.5733 0.01233
+Table 7: Stealthiness comparison on ImageNet.
 6 Conclusions and Future Work
-In this paper, we propose an imperceptible and effective backdoor attack against self-supervised models based on the optimized triggers that is disentangled to the augmented transformation in the self-supervised learning.
-Based on the evaluation across five different datasets and seven SSL algorithms, our attack is demonstrated to be both highly effective and stealthy.
+In this paper, we propose an imperceptible and effective backdoor attack against self-supervised models based on the optimized triggers that are disentangled in the augmented transformation in the SSL.
+Based on the evaluation across five different datasets and six SSL algorithms, our attack is demonstrated to be both highly effective and stealthy.
 It also effectively bypasses existing backdoor defenses.
 For future work, it would be beneficial to expand the scope of research to include various other domains of machine learning, such as NLP and audio processing.
-Appendix A More Attacking Cases of IMPERATIVE
-A.1 Generalization to Other Representative SSL Frameworks
-As elaborated on subsection 2.1, SSL frameworks for Contrastive Learning can be normally based on the following categories: negative examples, self-distillation, and clustering.
-To further evaluate the efficacy of IMPERATIVE, we test our method across the representative SSL algorithms in each category apart from the default SimCLR, i.e., negative examples (MoCo [22]), self-distillation (BYOL [17], SimSiam [6], and DINO [4]), and clustering (SwAV [3]).
-Table 7 illustrates the detailed parameters for different augmentation transforms used in the SSL algorithms, including Random resized crop, Random horizontal flip, Color jitter, Random gray scale, and Random polarization.
-The backbone feature extractor used here is ResNet18 [21].
-Other experiment settings are the same as those in Appendix D.6.3.
-Table 6 shows the evaluation results of applying IMPERATIVE to the encoders pre-trained utilizing various SSL algorithms on CIFAR-10.
-The SSL algorithms demonstrate high effectiveness in achieving successful backdoor attacks with ASR nearing or reaching perfection in many cases, with an average ASR of 99.54%.
-Additionally, the algorithms excel in maintaining the stealthiness of these attacks, evidenced by a high average SSIM of 0.95 and a high average PSNR value of 26.89, alongside a low LPIPS value of 0.011.
-These metrics across the SSL algorithms collectively indicate that while IMPERATIVE are highly effective at manipulating model behaviors, they do so in a way that remains largely undetectable through standard image quality assessments, which aligns with the conclusions drawn in section 5.
+Supplementary Material
+A1 Generalization to Other SSL Frameworks
+SSL frameworks for Contrastive Learning can be normally based on the following categories: negative examples, self-distillation, and clustering.
+To further evaluate the efficacy of INACTIVE, we test our method across the representative SSL algorithms in each category apart from the default SimCLR, i.e., negative examples (MoCo [20]), self-distillation (BYOL [16], SimSiam [7]), and clustering (SwAV [4]).
+Table A5 illustrates the detailed parameters for different augmentation transforms used in the SSL algorithms.
+Other experiment settings are the same as subsection A6.5.
+Experimental Results.
+Table A1 shows the evaluation results of applying INACTIVE to the encoders pre-trained utilizing various SSL algorithms on CIFAR-10.
+The SSL algorithms demonstrate high effectiveness in achieving successful backdoor attacks with ASR nearing or reaching perfection in many cases, with an average ASR of 99.62%.
+Additionally, the algorithms excel in maintaining the stealthiness of these attacks, evidenced by a high average SSIM of 0.95 and a high average PSNR value of 28.42, alongside a low LPIPS value of 0.0086.
+These metrics across the SSL algorithms collectively indicate that while INACTIVE is highly effective at manipulating model behaviors, they do so in a way that remains largely undetectable through standard image quality assessments.
 […]
@@
 SVHN 74.58 99.97 0.96 31.86 0.0030
-SimSiam [6] STL10 57.56 100.00 0.95 23.29 0.0230
+SimSiam [7] STL10 64.79 99.38 0.96 32.46 0.0019
 GTSRB 28.99 99.90 0.96 24.26 0.0190
 SVHN 52.24 100.00 0.96 26.04 0.0141
-BYOL [17] STL10 39.31 98.33 0.95 26.94 0.0070
-GTSRB 11.15 99.31 0.96 32.35 0.0020
-SVHN 42.76 99.74 0.96 31.88 0.0024
+BYOL [16] STL10 79.49 99.98 0.96 31.44 0.0025
+GTSRB 70.70 100.00 0.96 35.34 0.0011
+SVHN 42.76 99.74 0.96 32.12 0.0024
 SwAV [4] STL10 66.28 99.76 0.89 23.43 0.0258
@@
 SVHN 76.71 96.94 0.96 25.05 0.0096
-DINO [4] STL10 36.06 99.57 0.93 26.98 0.0141
-GTSRB 11.49 99.83 0.95 27.11 0.0069
-SVHN 37.66 99.99 0.93 25.18 0.0184
-Average 50.32 99.54 0.95 26.89 0.0110
-Table 6: Evaluations on other SSL algorithms, which are used to pre-train the encoders on CIFAR-10. IMPERATIVE keeps highly effective and stealthy across these algorithms.
-Parameter SimCLR/CLIP MOCO SimSiam BYOL SwAV DINO Description
-input_size 32/224 32 32 32 / / Size of the input image in pixels.
-cj_prob 0.8 0.8 0.8 0.8 0.8 0.8 Probability that color jitter is applied.
-cj_strength 1 0.5 1 1 0.5 0.5 Strength of the color jitter. cj_bright, cj_contrast, cj_sat, and cj_hue are multiplied by this value.
-cj_bright 0.4 0.4 0.4 0.4 0.8 0.8 How much to jitter brightness.
-cj_contrast 0.4 0.4 0.4 0.4 0.8 0.8 How much to jitter contrast.
-cj_sat 0.4 0.4 0.4 0.2 0.8 0.4 How much to jitter saturation.
-cj_hue 0.1 0.1 0.1 0.1 0.2 0.2 How much to jitter hue.
-min_scale / 0.2 0.2 0.08 / / Minimum size of the randomized crop relative to the input_size.
-random_gray_scale 0.2 0.2 0.2 0.2 0.2 0.2 Probability of conversion to grayscale.
-vf_prob 0 0 0 0 0 0 Probability that vertical flip is applied.
-hf_prob 0.5 0.5 0.5 0.5 0.5 0.5 Probability that horizontal flip is applied.
-rr_prob 0 0 0 0 0 0 Probability that random rotation is applied.
-crop_sizes / / / / [32,32] / Size of the input image in pixels for each crop category.
-crop_counts / / / / [2,2] / Number of crops for each crop category.
-crop_min_scales / / / / [0.14,0.14] / Min scales for each crop category.
-crop_max_scales / / / / [32,32] / Max_scales for each crop category.
-global_crop_size / / / / / 32 Crop size of the global views.
-global_crop_scale / / / / / [0.4, 1.0] Tuple of min and max scales relative to global_crop_size.
-local_crop_size / / / / / 96 Crop size of the local views.
-local_crop_scale / / / / / [0.05, 0.4] Tuple of min and max scales relative to local_crop_size.
-n_local_views / / / / / 6 Number of generated local views.
-solarization_prob / / / / / 0.2 Probability to apply solarization on the second global view.
-Table 7: Parameters of Augmentation Transforms in SSL Algorithms.
-A.2 Generalization to Multi-modal Models
-We utilize our IMPERATIVE to integrate a backdoor into the image encoder of the multi-modal model CLIP [58], which is comprised of both an image and a text encoder, pre-trained on 400 million internet-sourced (image, text) pairs.
+Average 64.72 99.62 0.95 28.42 0.0086
+Table A1: Evaluations on other SSL algorithms, which are used to pre-train the encoders on CIFAR-10. INACTIVE keeps highly effective and stealthy across these algorithms.
+[…]
+Ours 70.6 68.98 99.95 0.98 30.11 0.0179
+BadEncoder 70.27 99.99 0.28 9.98 0.6105
+Table A2: Attack results on the multi-shot classification of CLIP. INACTIVE attains comparable high BA and ASR to those of BadEncoder [27], yet it offers significantly enhanced stealthiness.
+A2 Generalization to Multi-modal Models
+We utilize our INACTIVE to integrate a backdoor into the image encoder of the multi-modal model CLIP [53], which is comprised of both an image and a text encoder, pre-trained on 400 million internet-sourced (image, text) pairs.
 This backdoored image encoder is then employed to construct a multi-shot downstream classifier.
-Due to the unavailability of CLIP’s original pre-training dataset, we utilize the training images from CIFAR-10, resized to 224×224×3, as our shadow dataset.
-We define $\alpha=0.1$ and $\beta=20$ in Algorithm 1, and $\mu=0.1$ in Algorithm 2 to normalize the loss terms across similar scales.
-Additionally, Algorithm 1 and Algorithm 2 are configured with a batch size of 16, and learning rates are set at $10^{-6}$ and 0.005, respectively, for a duration of 100 epochs.
-SVHN serves as the downstream dataset with the target label “one”.
-We employ the same reference inputs as described in [29].
-All other experimental settings remain consistent with those outlined in Appendix D.6.3.
-Table 8 presents the attack results of multi-shot classification using CLIP.
-IMPERATIVE achieves a similar high BA and ASR as BadEncoder [29] does, but it maintains much more stealthiness with high SSIM&PSNR and low LPIPIS values than BadEncoder which uses a white patch as the backdoor trigger.
-Methods CA(%) BA(%) ASR(%) SSIM PSNR LPIPIS
-Ours 70.6 68.98 99.95 0.98 30.11 0.0179
-BadEncoder [29] 70.27 99.99 / / /
-Table 8: Attack results on the multi-shot classification of CLIP. IMPERATIVE attains comparable high BA and ASR to those of BadEncoder [29], yet it offers significantly enhanced stealthiness.
-Appendix B Ablation Study
-As discussed in subsection 4.4, we utilize $\mathcal{L}_{disentangle}$ to enlarge the gap between the backdoored image distribution and the augmented image distribution.
-We use $\mathcal{L}_{alignment}$ to promote feature alignment between backdoored images and reference images.
-We use $\mathcal{L}_{stealthy}$ to make the backdoor trigger stealthy and natural.
-In Algorithm 1, we initialize a backdoor injector using a pre-trained one by Algorithm 2.
+Experimental Results.
+Table A2 presents the attack results of multi-shot classification using CLIP.
+INACTIVE achieves a similar high BA and ASR as BadEncoder [27] does, but it maintains much more stealthiness with high SSIM&PSNR and low LPIPIS values than BadEncoder which uses a white patch as the backdoor trigger.
+A3 Ablation Study
+We utilize $\mathcal{L}_{\text{disentangle}}$ to enlarge the gap between the backdoored image distribution and the augmented image distribution.
+We use $\mathcal{L}_{\text{alignment}}$ to promote feature alignment between backdoored images and reference images.
+We use $\mathcal{L}_{\text{stealthy}}$ to make the backdoor trigger stealthy and natural.
+In Alg.
+1, we initialize a backdoor injector using a pre-trained one by Alg.
+2.
 We use STL10 as the dataset for pre-training encoders.
-In this section, we demonstrate the effects of each important component in subsection 4.4 by ablating them respectively.
-As reported in Table 9, the ASRs of our method decrease across all target downstream datasets when ablating each component respectively.
-As reported in Table 10, the average SSIM and PSNR drop dramatically and LPIPS increases a lot across all datasets when ablating $\mathcal{L}_{stealthy}$.
-For example, in the ablation case of the target downstream dataset CIFAR10, the ASR decreases from 99.58% to 67.84%, 10.22%, 75.27% when ablating $\mathcal{L}_{disentangle}$, $\mathcal{L}_{alignment}$ and backdoor injector initialization respectively.
+In this section, we demonstrate the effects of each important component by ablating them respectively.
+As reported in Table A3, the ASRs of our method decrease across all target downstream datasets when ablating each component respectively.
+As reported in Table A4, the average SSIM and PSNR drop dramatically and LPIPS increases a lot across all datasets when ablating $\mathcal{L}_{\text{stealthy}}$.
+For example, in the ablation case of the target downstream dataset CIFAR10, the ASR decreases from 99.58% to 67.84%, 10.22%, 75.27% when ablating $\mathcal{L}_{\text{disentangle}}$, $\mathcal{L}_{\text{alignment}}$ and backdoor injector initialization respectively.
 This distinctly shows that the absence of key components significantly compromises the effectiveness of our backdoor attack methodology.
-Moreover, when ablating $\mathcal{L}_{stealthy}$, the SSIM drops from 0.99 to 0.1.
+Moreover, when ablating $\mathcal{L}_{\text{stealthy}}$, the SSIM drops from 0.99 to 0.1.
 The PSNR drops from 41.80 to 5.65.
-The LPIPS increases from 0.0002 to 0.55.
-This indicates that the backdoored images without the $\mathcal{L}_{stealthy}$ are more easily detectable and differ a lot from the original images.
+The LPIPS increases from 0.0002 to 0.55 indicating that the backdoored images without the $\mathcal{L}_{\text{stealthy}}$ are more easily detectable and differ a lot from the original images.
 Pre-training Backdoor injector $\mathcal{L}_{\text{disentangle}}$ $\mathcal{L}_{\text{alignment}}$ $\mathcal{L}_{\text{stealthy }}$ Target Downstream Dataset
@@
 87.11 99.58 75.82 97.97 58.62 99.76
-Table 9: ASR(%), BA(%) in ablation studies.
-Ablating the $\mathcal{L}_{\text{alignment}}$ means ablating the one in $\mathcal{L}_{\text{injector}}$ instead of $\mathcal{L}_{encoder}$, and ablating the pre-trained backdoor injector means randomly initializing it.
+Table A3: ASR(%), BA(%) in ablation studies.
+Ablating the $\mathcal{L}_{\text{alignment}}$ means ablating the one in $\mathcal{L}_{\text{injector}}$ instead of $\mathcal{L}_{\text{encoder}}$, and ablating the pre-trained backdoor injector means randomly initializing it.
 Method Target Downstream Dataset
@@
 […]
-w/o $\mathcal{L}_{stealthy}$ 0.10 5.65 0.5500 0.02 5.20 0.4400 0.10 4.85 0.52
+w/o $\mathcal{L}_{\text{stealthy}}$ 0.10 5.65 0.5500 0.02 5.20 0.4400 0.10 4.85 0.52
 Ours 0.99 41.80 0.0002 0.99 43.58 0.0001 0.96 25.65 0.01
-Table 10: Results of ablating $\mathcal{L}_{stealthy}$.
-Images backdoored without the $\mathcal{L}_{stealthy}$ are more readily identifiable and exhibit significant deviations from the original images.
-Appendix C More Evaluations
-C.1 More Robustness Tests
-C.1.1 Beatrix
+Table A4: Results of ablating $\mathcal{L}_{\text{stealthy}}$.
+Images backdoored without the $\mathcal{L}_{\text{stealthy}}$ are more readily identifiable and exhibit significant deviations from the original images.
+Parameter SimCLR/CLIP MOCO SimSiam BYOL SwAV Description
+input_size 32/224 32 32 32 / Size of the input image in pixels.
+cj_prob 0.8 0.8 0.8 0.8 0.8 Probability that color jitter is applied.
+cj_strength 1 0.5 1 1 0.5 Strength of the color jitter. cj_bright, cj_contrast, cj_sat, and cj_hue are multiplied by this value.
+cj_bright 0.4 0.4 0.4 0.4 0.8 How much to jitter brightness.
+cj_contrast 0.4 0.4 0.4 0.4 0.8 How much to jitter contrast.
+cj_sat 0.4 0.4 0.4 0.2 0.8 How much to jitter saturation.
+cj_hue 0.1 0.1 0.1 0.1 0.2 How much to jitter hue.
+min_scale / 0.2 0.2 0.08 / Minimum size of the randomized crop relative to the input_size.
+random_gray_scale 0.2 0.2 0.2 0.2 0.2 Probability of conversion to grayscale.
+vf_prob 0 0 0 0 0 Probability that vertical flip is applied.
+hf_prob 0.5 0.5 0.5 0.5 0.5 Probability that horizontal flip is applied.
+rr_prob 0 0 0 0 0 Probability that random rotation is applied.
+crop_sizes / / / / [32,32] Size of the input image in pixels for each crop category.
+crop_counts / / / / [2,2] Number of crops for each crop category.
+crop_min_scales / / / / [0.14,0.14] Min scales for each crop category.
+crop_max_scales / / / / [32,32] Max scales for each crop category.
+Table A5: Parameters of Augmentation Transforms in SSL Algorithms.
+A4 Proof
+We aim to prove the following statement:
+Theorem 3.1.
+Given a perfectly-trained encoder $\mathcal{F}_{\theta}$ based on the augmentations sampled from the predefined augmentation space $S_{A}$, it is impossible to inject a backdoor with a trigger function $\mathcal{T}\in S_{A}$.
+Proof.
+The encoder $\mathcal{F}_{\theta}$ is said to be perfectly-trained if it achieves maximal similarity on any pair of augmented versions of the same training sample $x$, for all augmentations in the augmentation space $S_{A}$.
+Formally, this means:
+$s(\mathcal{F}_{\theta}(A_{1}(x)),\mathcal{F}_{\theta}(A_{2}(x)))=1,\quad\forall x\in\mathcal{X},\,\forall A_{1},A_{2}\in S_{A},$
+where $s(\cdot,\cdot)$ denotes the similarity measure, which achieves its maximum value when the two inputs are identical in the feature space.
+As a result, for all $x\in\mathcal{X}$ and any $A_{1},A_{2}\in S_{A}$, the representations produced by $\mathcal{F}_{\theta}$ are identical:
+$\mathcal{F}_{\theta}(A_{1}(x))=\mathcal{F}_{\theta}(A_{2}(x)).$
+A backdoor injection introduces a trigger function $\mathcal{T}$ such that for any input $x$, the representation produced by $\mathcal{F}_{\theta}$ after applying the trigger, $\mathcal{F}_{\theta}(\mathcal{T}(x))$, maps to a target representation that is distinct from those of legitimate augmentations.
+In this case, we assume $\mathcal{T}\in S_{A}$, meaning the trigger function is an augmentation within the predefined augmentation space.
+Since $\mathcal{T}\in S_{A}$, by the definition of the perfectly-trained encoder, the following holds for all $x\in\mathcal{X}$ and any $A_{1},A_{2}\in S_{A}$:
+$\mathcal{F}_{\theta}(A_{1}(x))=\mathcal{F}_{\theta}(\mathcal{T}(x))=\mathcal{F}_{\theta}(A_{2}(x)).$
+This implies that the feature representation of the input $x$ after applying the trigger $\mathcal{T}$ is indistinguishable from those produced by any other augmentations in $S_{A}$.
+As such, the trigger $\mathcal{T}$ fails to produce a unique or distinct representation in the feature space.
+For a backdoor to be successfully injected, the trigger function $\mathcal{T}$ must produce a representation $\mathcal{F}_{\theta}(\mathcal{T}(x))$ that is distinct from those produced by other augmentations in $S_{A}$.
+However, this contradicts the property of the perfectly-trained encoder:
+$\mathcal{F}_{\theta}(A_{1}(x))=\mathcal{F}_{\theta}(\mathcal{T}(x)),\quad\forall A_{1}\in S_{A}.$
+Thus, $\mathcal{T}$ cannot create a distinct feature representation and, therefore, cannot function as a backdoor.
+We have shown that for a perfectly-trained encoder $\mathcal{F}_{\theta}$, trained on augmentations from $S_{A}$, any trigger function $\mathcal{T}\in S_{A}$ fails to produce a distinct representation necessary for a backdoor injection.
+Therefore, it is impossible to inject a backdoor with a trigger function $\mathcal{T}\in S_{A}$.
+A5 More Evaluations
+A5.1 More Robustness Tests
+Beatrix.
 Further examination of Beatrix’s detection performance reveals that it determines whether a sample is poisoned based on the deviation in the feature matrix within the embedding space [44].
 A sample is deemed poisoned if the deviation exceeds a specific threshold.
-Figure 4 displays the distribution of deviation values.
+Figure A1 displays the distribution of deviation values.
 Notably, poisoned samples by BadEncoder show significantly higher deviation than clean samples, enabling Beatrix to achieve high detection accuracy.
-Conversely, the deviation distributions between clean inputs and poisoned samples by IMPERATIVE overlap considerably, making them difficult to differentiate.
-Figure 4: Beatrix’s deviation distribution [48] for clean and backdoored data, where blue bars represent clean samples and orange bars indicate poisoned ones.
+Conversely, the deviation distributions between clean inputs and poisoned samples by INACTIVE overlap considerably, making them difficult to differentiate.
+Figure A1: Beatrix’s deviation distribution [44] for clean and backdoored data, where blue bars represent clean samples and orange bars indicate poisoned ones.
 Poisoned samples from BadEncoder display markedly higher deviations than clean samples, allowing Beatrix to detect them effectively.
-However, the deviation distributions for clean inputs and poisoned samples from IMPERATIVE significantly overlap, complicating their differentiation.
-Figure 5: Resilience to Grad-CAM [64]. The resemblance observed in these heatmaps indicates that IMPERATIVE is capable of resisting defenses that rely on Grad-CAM.
-C.1.2 Grad-CAM
-We assess IMPERATIVE’s resilience to Grad-CAM [64], which generates a heatmap for a given model and input sample, where the heat intensity of each pixel reflects its significance in the model’s final prediction.
+However, the deviation distributions for clean inputs and poisoned samples from INACTIVE significantly overlap, complicating their differentiation.
+Neural Cleanse.
+We then test the resilience of our method against a well-known reverse engineering defense, namely Neural Cleanse (NC) [72].
+Specifically, it utilizes the anomaly index to measure the deviation of the reconstructed triggers by their sizes, labeling models with an anomaly index exceeding two as Trojan-infected.
+Since NC is specifically designed for classifiers and cannot be directly used on image encoders, NC is employed to identify backdoors in a compromised downstream classifier.
+Moreover, the testing dataset from a targeted downstream dataset serves as the clean dataset.
+Experimental outcomes, as illustrated in Table A6, indicate that NC is unable to detect the Trojan model created by our approach.
+While effective in identifying patch-based Trojans [40], this method presumes a uniform trigger pattern at the pixel level across different samples.
+Our method circumvents this by employing input-dependent triggers, meaning the pixel-level Trojan alterations vary across samples.
+Pre-trained Dataset Downstream Dataset Anomaly Index  [72]
+CIFAR10 STL10 1.17
+SVHN 1.42
+GTSRB 1.61
+STL10 CIFAR10 0.88
+SVHN 1.23
+GTSRB 1.43
+Table A6: Evaluation results of neural cleanse [72]. A model is judged as backdoored if its anomaly index >2, so our attack cannot be detected by it.
+Figure A2: Resilience to Grad-CAM [58]. The resemblance observed in these heatmaps indicates that INACTIVE is capable of resisting defenses that rely on Grad-CAM.
+Grad-CAM.
+We assess INACTIVE’s resilience to Grad-CAM [58], which generates a heatmap for a given model and input sample, where the heat intensity of each pixel reflects its significance in the model’s final prediction.
 Grad-CAM is effective for identifying smaller Trojans [41, 17], as these Trojans tend to produce high heat values in compact trigger zones, leading to abnormal heatmaps.
 Nevertheless, our backdoor transformation function alters the entire image, rendering Grad-CAM ineffective in detecting it.
-Figure 5 illustrates the comparison of heatmaps from both clean images and backdoored images created by our method.
-The similarity in these heatmaps suggests that IMPERATIVE can withstand defenses based on Grad-CAM.
-C.1.3 Various Noises
-Next, we are going to introduce the noises, i.e., JPEG compression, Salt&Pepper noise, and Poisson noise, applied in the robustness test.
+Figure A2 illustrates the comparison of heatmaps from both clean images and backdoored images created by our method.
+The similarity in these heatmaps suggests that INACTIVE can withstand defenses based on Grad-CAM.
+Various Noises. Next, we are going to introduce the noises, i.e., JPEG compression, Salt&Pepper noise, and Poisson noise, applied in the robustness test.
   • JPEG Compression: JPEG is a prevalent image format.
@@
 It typically ranges from 0 to 1, but practical applications often use a much smaller range (e.g., 0.01 to 0.2) to avoid overwhelming the image with noise.
-Figure 6: Resilience to JPEG compression, Salt&Pepper noise, Poisson noise of different densities.
+Figure A3: Resilience to JPEG compression, Salt&Pepper noise, Poisson noise of different densities.
 Evaluating the robustness of our method across these noises helps ensure its performance is reliable even when image quality is compromised when encountering real-world image issues.
-We illustrate the evaluation results in Figure 6, where the pre-trained dataset is CIFAR10 and the downstream dataset is STL10.
-The results prove that our method, which utilizes $\mathcal{L}_{disentangle}$, maintains high ASRs in most cases, and the ASR is significantly higher than the cases ablating $\mathcal{L}_{disentangle}$ in our method.
-For example, under different intensities of Poisson noise, the ASR of our method maintains nearly 100%, while the ASR of the cases ablating $\mathcal{L}_{disentangle}$ in our method is unstable, ranging from 12% to 61%.
-The outcome indicates that the $\mathcal{L}_{disentangle}$ plays a prominent role in promoting the robustness of our method.
-The reason may be that the $\mathcal{L}_{disentangle}$ encourages the model to learn features that remain invariant under the transformations.
+We illustrate the evaluation results in Figure A3, where the pre-trained dataset is CIFAR10 and the downstream dataset is STL10.
+The results prove that our method, which utilizes $\mathcal{L}_{\text{disentangle}}$, maintains high ASRs in most cases, and the ASR is significantly higher than the cases ablating $\mathcal{L}_{\text{disentangle}}$ in our method.
+For example, under different intensities of Poisson noise, the ASR of our method maintains nearly 100%, while the ASR of the cases ablating $\mathcal{L}_{\text{disentangle}}$ in our method is unstable, ranging from 12% to 61%.
+The outcome indicates that the $\mathcal{L}_{\text{disentangle}}$ plays a prominent role in promoting the robustness of our method.
+The reason may be that the $\mathcal{L}_{\text{disentangle}}$ encourages the model to learn features that remain invariant under the transformations.
 Namely, the model can still recognize the embedded triggers even when various noises or other image perturbations are applied.
 This aspect of feature preservation is key to ensuring that the backdoor attack remains effective regardless of the image transformations or certain noises applied.
-C.1.4 STRIP
-Figure 7: Resilient to STRIP [16]. IMPERATIVE can withstand the STRIP runtime defense.
+Figure A4: Resilient to STRIP [14]. INACTIVE can withstand the STRIP runtime defense.
+STRIP.
 STRIP [14] scrutinizes a suspect sample by overlaying diverse image patterns onto it and monitoring the uniformity of the model’s predictions for these altered inputs.
 A low entropy score, suggesting consistent predictions across these perturbed samples, would lead STRIP to classify the sample as Trojan-infected.
-The experimental outcomes depicted in Figure 7 reveal that the entropy values for both clean and Trojan-compromised models produced through our method overlap considerably, signifying that our attack can withstand the STRIP runtime defense.
-IMPERATIVE manages to evade STRIP as the superimposition process destroys the attack’s trigger, thereby causing both the Trojan-induced and clean samples to exhibit significant alterations in prediction when superimposed, aligning them with what is observed in clean cases.
-C.2 Sensitivity Evaluation
-Figure 8: Sensitivity evaluation of hyperparameters. IMPERATIVE performs consistently over these parameter ranges.
-Figure 8 illustrates the sensitivity evaluation of the hyperparameters alpha and beta, showcasing that the model’s performance exhibits relative insensitivity to parameter variations.
+The experimental outcomes depicted in Figure A4 reveal that the entropy values for clean and Trojan-compromised models produced through our method overlap considerably, signifying that our attack can withstand the STRIP runtime defense.
+INACTIVE manages to evade STRIP as the superimposition process destroys the attack’s trigger, thereby causing both the Trojan-induced and clean samples to exhibit significant alterations in prediction when superimposed, aligning them with what is observed in clean cases.
+Adaptive Defense.
+Our threat model assumes the attacker controls pre-downstream training, while the defender can only intervene during downstream training and inference.
+Since our trigger injection operates in the HSV color space, we introduce Salt and Pepper Noise and Poisson Noise perturbations in the HSV color space of both downstream training and inference phase, simulating potential adaptive defense strategies a defender might use to disrupt the attack.
+We use CIFAR10 as the pre-training dataset and GTSRB, STL10, and SVHN as downstream datasets.
+Experimental results show INACTIVE’s ASR remains unaffected, demonstrating its robustness.
+This resilience stems from our disentanglement loss, which increases the distributional gap between backdoor and augmented samples.
+This design inherently enhances the robustness of the trigger, as the larger distributional gap in the color space ensures that the trigger remains distinguishable even when noise is introduced to the samples.
+Consequently, the trigger is less affected by noisy samples and retains its effectiveness in activating the backdoor.
+Downstream dataset Salt and Pepper on HSV Poisson Noise on HSV
+BA ASR BA ASR
+GTSRB 81.96% 100% 81.81% 100%
+STL10 73.81% 99.38% 73.65% 99.41%
+SVHN 60.17% 95.34% 60.65% 91.11%
+Table A7: Adaptive defense evaluation results pre-trained on CIFAR10.
+A5.2 Sensitivity Evaluation
+Figure A5 illustrates the sensitivity evaluation of the hyper-parameters alpha and beta, showcasing that the model’s performance exhibits relative insensitivity to parameter variations.
 Specifically, for alpha, both the Attack Success Rate (ASR) and Backdoored Accuracy (BA) remain relatively constant across a range of alpha values from 0.05 to 0.30, suggesting that once an optimal threshold is reached, the model’s performance is stable despite further adjustments to alpha.
@@
 The horizontal axis in both graphs represents the incremental values of alpha and beta, demonstrating the model’s consistent performance over these parameter ranges.
-C.3 Results Analysis
-Figure 9: Visualization of feature space segregation over epochs demonstrates the efficacy of our backdoor trigger in creating a distinguishable separation between backdoor and clean features, unlike WaNet [51] where the features remain intermingled, highlighting our method’s utility in securing encoders against backdoor threats.
-Figure 9 presents a visual analysis of the feature shifts induced by backdoor injection across different epochs for our proposed method compared to the WaNet [51].
+Figure A5: Sensitivity evaluation of hyper-parameters. INACTIVE performs consistently over these parameter ranges.
+A5.3 Results Analysis
+Figure A6: Visualization of feature space segregation over epochs demonstrates the efficacy of our backdoor trigger in creating a distinguishable separation between backdoor and clean features, unlike WaNet [47] where the features remain intermingled, highlighting our method’s utility in securing encoders against backdoor threats.
+Figure A6 presents a visual analysis of the feature shifts induced by backdoor injection across different epochs for our proposed method compared to the WaNet [47].
 Throughout the backdoor injection process, it becomes evident that our method’s trigger effectively separates the backdoor features from the clean features.
@@
 It introduces a discernible feature shift that an encoder can leverage to recognize backdoored inputs, validating the practical utility of our approach in enhancing model security against backdoor attacks.
-Appendix D Implementation Details
-The following sections provide a comprehensive breakdown of the process of implementing the attack, which is methodically organized into three distinct phases: pre-training the encoder (Appendix D.1), pre-training the backdoor injector (Appendix D.2), injecting the backdoor (Appendix D.3), and training the downstream classifiers (Appendix D.4).
-Additionally, we explain the HSV&HSL Color Spaces and Wasserstein Distance (Appendix D.5) used in the method in detail.
-Moreover, we provide more details on the datasets (Appendix D.6.1), evaluation metrics (Appendix D.6.2), and experimental setup details (Appendix D.6.3, Appendix D.6.4).
-D.1 Pre-training Image Encoders
+[…]
+Epoch 10 20 50 100
+[…]
+[…]
+Table A8: KL divergence values across different epochs.
+A5.4 Stealthiness Evaluation
+A5.4.1 Algorithmic Metrics
+Table A9 and Table A10 show the detailed experimental results of the stealthy metrics across various backdoor attack methods and datasets.
+[…]
+Badencoder [27] /DRUPE [68] CIFAR10 STL10 0.8375 14.3745 0.03894 0.7788 63.7301
+SVHN 0.8380 15.0739 0.11092 0.8518 33.1811
+GTSRB 0.8309 12.8845 0.08094 0.8295 63.1784
+STL10 CIFAR10 0.8455 14.9113 0.03696 0.8598 13.8215
+SVHN 0.8407 15.0849 0.11110 0.8165 64.5712
+GTSRB 0.8308 12.8965 0.08597 0.8306 63.2014
+Average 0.8372 14.2043 0.07747 0.8278 50.2806
+CTRL [34] CIFAR10 STL10 0.9415 32.4195 0.00017 0.9115 47.7871
+SVHN 0.8695 32.3533 0.00065 0.7032 157.6731
+GTSRB 0.8965 32.4565 0.00021 0.9787 7.9540
+STL10 CIFAR10 0.9356 32.4603 0.00019 0.9827 11.9283
+SVHN 0.8721 32.4198 0.00062 0.9347 59.4134
+GTSRB 0.8945 32.5226 0.00024 0.9773 8.8889
+Average 0.9016 32.4387 0.00035 0.9147 48.9408
+WaNet [47] CIFAR10 STL10 0.7378 14.3472 0.04657 0.6611 97.1471
+SVHN 0.7941 15.2503 0.10023 0.5927 129.6770
+GTSRB 0.7792 13.1140 0.07615 0.7320 67.4510
+STL10 CIFAR10 0.7728 14.9430 0.03996 0.6510 96.0399
+SVHN 0.8026 15.4044 0.09544 0.5924 129.8306
+GTSRB 0.7687 13.1645 0.08358 0.7314 67.7673
+Average 0.7759 14.3706 0.07366 0.6601 97.9855
+Ins-Kelvin [42] CIFAR10 STL10 0.4427 15.7024 0.11110 0.7011 107.4607
+SVHN 0.4213 15.4081 0.20070 0.5905 121.0396
+GTSRB 0.6224 17.4669 0.10820 0.7393 60.8467
+STL10 CIFAR10 0.4624 15.6372 0.10970 0.6582 85.2866
+SVHN 0.4358 15.4436 0.19840 0.6342 130.1333
+GTSRB 0.5942 17.5105 0.11640 0.7422 62.0004
+Average 0.4965 16.1948 0.14075 0.6776 94.4612
+Ins-Xpro2 [42] CIFAR10 STL10 0.5672 17.8706 0.03451 0.7806 66.9640
+SVHN 0.5324 17.4862 0.06699 0.7698 27.4974
+GTSRB 0.6947 18.3951 0.03151 0.8997 10.7916
+STL10 CIFAR10 0.5354 17.9291 0.03806 0.8063 19.0678
+SVHN 0.5057 17.4956 0.07152 0.7505 45.3211
+GTSRB 0.6541 18.4070 0.03632 0.8932 11.9285
+Average 0.5816 17.7640 0.04615 0.8167 30.2617
+POIENC [39] CIFAR10 CIFAR10 0.1214 11.2787 0.15867 0.5967 172.2200
+BLTO [63] CIFAR10 CIFAR10 0.8417 29.6756 0.00941 0.9501 36.3848
+SSLBKD [56] CIFAR10 CIFAR10 0.8737 16.2414 0.09640 0.8913 118.3200
+Ours CIFAR10 STL10 0.9405 24.0804 0.02311 0.9161 45.8759
+SVHN 0.9807 46.0749 0.00040 0.9976 1.8810
+GTSRB 0.9687 37.4394 0.00338 0.9944 1.2025
+STL10 CIFAR10 0.9928 46.5797 0.00019 0.9981 0.5018
+SVHN 0.9847 45.7410 0.00031 0.9464 30.0092
+GTSRB 0.9905 46.5149 0.00011 0.9977 0.2157
+Average 0.9763 41.0717 0.0046 0.9751 13.2810
+Table A9: Stealthiness Metrics across Methods, Pre-training and Downstream Datasets.
+Model Pre-training Dataset Downstream Dataset SSIM PSNR(dB) LPIPS
+ISSBA [36] ImageNet STL10 0.7836 30.3333 0.05615
+GTSRB 0.7292 31.7344 0.11651
+SVHN 0.6860 31.9810 0.20006
+Average 0.7329 31.3496 0.12424
+Ours ImageNet STL10 0.9900 38.3300 0.00400
+GTSRB 0.9900 32.1400 0.01700
+SVHN 0.9800 33.2500 0.01600
+Average 0.9867 34.5733 0.01233
+Table A10: Stealthiness Metrics across Methods on ImageNet.
+A5.4.2 Human Inspection
+We conducted a human inspection study [47] to evaluate the stealthiness of various backdoor attacks.
+We introduce the detailed settings in section A6.
+The results in Table A11 indicate that traditional methods like SSLBKD, BadEncoder/DRUPE, and POIENC have relatively low fooling rates, with SSLBKD at just 4.9% overall, suggesting they are easier for humans to detect.
+In contrast, our method achieves a higher fooling rate of 49.5% overall, indicating much greater stealth.
+However, as shown in Figure A9, subtle artifacts, such as slight color changes, can still be detected by humans.
+Images SSLBKD CTRL BadEncoder /DRUPE POIENC Ins-Kelvin BLTO ISSBA Ours
+Backdoor 2.4 31.2 4.0 9.6 12.8 22.4 24.8 50.6
+Clean 7.4 27.2 8.8 16.8 10.4 27.6 26.4 48.4
+Both 4.9 29.2 6.4 13.2 11.6 25.0 25.6 49.5
+Table A11: Success fooling $\uparrow$ (%) of different methods. Our method achieves the greatest stealth with the highest overall success fooling rates.
+A6 Implementation Details
+The following sections provide a comprehensive breakdown of the process of implementing the attack, which is methodically organized into three distinct phases: pre-training the encoder, injecting the backdoor, and training the downstream classifiers.
+Additionally, we explain the HSV&HSL Color Spaces and Wasserstein Distance used in the method in detail.
+Moreover, we provide more details on the datasets and experimental setup details.
+A6.1 Pre-training Image Encoders
 The initial step involves pre-training an image encoder using a specific dataset, hereinafter referred to as the pre-training dataset.
@@
 Meanwhile, for STL10, the pre-training also incorporates its additional unlabeled images.
-D.2 Pre-training Backdoor Injector
-To enhance the optimization speed and increase the effectiveness of the backdoor attack, we use the method outlined in Algorithm 2 to pre-train a backdoor injector, which serves as the initial setup for the injector in Algorithm 1.
-We use the U-Net architecture [59] for the backdoor injector.
-The detailed architecture is shown in the Table 11.
-Algorithm 2 Pre-training backdoor injector
-Input: Shadow dataset $\mathcal{D}_{s}$       Output: Pre-trained backdoor injector $\mathcal{I}_{\phi}$
-1:function Pre-training injector($\mathcal{D}_{s}$)
-2:     $\hat{\mathcal{D}_{s}}\leftarrow$ Augment samples in $\mathcal{D}_{s}$; $\mathcal{I}_{\phi}\leftarrow\text{Random initialization}$
-3:     for $\text{iter}=0$ to max_epochs do
-4:         ${\mathcal{D}_{s}}^{{}^{\prime}}\leftarrow\mathcal{I}_{\phi}(\mathcal{D}_{s})$
-5:         $\mathcal{L}_{\text{disentangle}}\leftarrow$ distribution difference between backdoor images $\boldsymbol{x}^{\prime}$ and augmented images $\tilde{\boldsymbol{x}}$, $\forall\boldsymbol{x}^{\prime}\in{\mathcal{D}_{s}}^{{}^{\prime}},\forall\tilde{\boldsymbol{x}}\in\hat{\mathcal{D}_{s}}$triangleright Equation 2
-6:         $\mathcal{L}_{\text{stealthy}}\leftarrow$ distance between backdoor image $\boldsymbol{x}^{\prime}$ and clean image $\boldsymbol{x}$, $\forall\boldsymbol{x}^{\prime}\in{\mathcal{D}_{s}}^{{}^{\prime}},\forall\boldsymbol{x}\in\mathcal{D}_{s}$
-7:         $\triangleright$ Equation 4
-8:         $\mathcal{L}_{\text{ours}}=\mathcal{L}_{\text{stealthy}}+\mu\cdot\mathcal{L}_{\text{disentangle}}$
-9:         $\phi_{\mathcal{I}}=\phi_{\mathcal{I}}-lr\cdot\frac{\partial\mathcal{L}_{\text{ours}}}{\partial\phi_{\mathcal{I}}}$
 Module Layer Type Kernel Size Stride Padding Channel I/O Normalization Activation Input
@@
 Conv_1x1 Conv2d 1x1 1 - 8/3 - - Up_conv2
-Table 11: The detailed network structure of backdoor injector.
-D.3 Backdoor Injection
+Table A12: The detailed network structure of backdoor injector.
+A6.2 Backdoor Injection
 The subsequent phase introduces a backdoor into the pre-trained encoder.
@@
 Additionally, we employ the same reference samples as those used in BadEncoder [27].
-The backdoor is injected using a designated algorithm, which leverages a combination of $\mathcal{L}_{consistency}$, $\mathcal{L}_{utility}$, and $\mathcal{L}_{alignment}$.
+The backdoor is injected using a designated algorithm, which leverages a combination of $\mathcal{L}_{\text{consistency}}$, $\mathcal{L}_{\text{utility}}$, and $\mathcal{L}_{\text{alignment}}$.
 The notations are the same as the main paper.
 Consistency Loss. The consistency loss decreases when the feature vectors generated by the backdoored image encoder and the clean image encoder for the reference inputs are more alike.
-$\mathcal{L}_{consistency}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\sum_{i=1}^{t}r_{i}}$ (6)
+$\mathcal{L}_{\text{consistency}}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\sum_{i=1}^{t}r_{i}}$ (A1)
 Utility Loss. Utility loss diminishes when the backdoored image encoder and the clean image encoder yield more closely matching feature vectors for a clean input within the shadow dataset.
-$\mathcal{L}_{utility}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}(\boldsymbol{x}),\mathcal{F}_{\theta}(\boldsymbol{x})\right)$ (7)
-D.4 Training Downstream Classifiers
+$\mathcal{L}_{\text{utility}}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}(\boldsymbol{x}),\mathcal{F}_{\theta}(\boldsymbol{x})\right)$ (A2)
+A6.3 Training Downstream Classifiers
 Finally, using the pre-trained image encoder, we proceed to train classifiers for three separate datasets, hereafter referred to as downstream datasets.
@@
 The training of the classifier is conducted over 500 epochs using the cross-entropy loss function and the Adam optimizer, with a set learning rate of 0.0001.
-D.5 Technical Background
+A6.4 Technical Background
 HSV Color Space.
+HSV and HSL color spaces align more closely with human color perception, unlike RGB’s additive mixing.
+They allow for intuitive adjustments of tint, shade, and tone through single-dimension modifications.
+They have wide applications in the computer vision domain [38, 12], as detailed in subsection A6.4.
 The HSV color model describes color based on three primary attributes: hue, saturation, and value.
@@
 Value (V), often referred to as brightness, measures the lightness or darkness of the color [61].
-HSV color space can be converted from RGB color space through Equation 8, Equation 9, Equation 10, Equation 11 [63].
+HSV color space can be converted from RGB color space through Equation A3, Equation A4, Equation A5, Equation A6 [57].
 HSL Color Space.
 The HSL color model is defined by three cylindrical coordinates: hue (H), saturation (S), and lightness (L), which together capture the subtleties of color.
-HSL color space can be converted from RGB color space through Equation 8, Equation 9, Equation 12, Equation 13 [63].
-$\mathrm{R}^{\prime}=\mathrm{R}/255;$ (8)
+HSL color space can be converted from RGB color space through Equation A3, Equation A4, Equation A7, Equation A8 [57].
+$\mathrm{R}^{\prime}=\mathrm{R}/255;$ (A3)
 $\mathrm{G}^{\prime}=\mathrm{G}/255;$
-$\mathrm{B}^{\prime}=\mathrm{B}/255$
-$\mathrm{C}_{\max}=\operatorname{MAX}\left(\mathrm{R}^{\prime},\mathrm{G}^{\prime},\mathrm{B}^{\prime}\right)$
-$\mathrm{C}_{\min}=\operatorname{MIN}\left(\mathrm{R}^{\prime},\mathrm{G}^{\prime},\mathrm{B}^{\prime}\right)$
-$\Delta=\mathrm{C}_{\max}-\mathrm{C}_{\min}$
-$\mathrm{H}=\left\{\begin{array}[]{l}60^{\circ}\times\left(\frac{\mathrm{G}^{\prime}-\mathrm{B}^{\prime}}{\Delta}\bmod 6\right),\mathrm{C}_{\max}=\mathrm{R}^{\prime}\\ 60^{\circ}\times\left(\frac{\mathrm{B}^{\prime}-\mathrm{R}^{\prime}}{\Delta}+2\right),\mathrm{C}_{\max}=\mathrm{G}^{\prime}\\ 60^{\circ}\times\left(\frac{\mathrm{R}^{\prime}-\mathrm{G}^{\prime}}{\Delta}+4\right),\mathrm{C}_{\max}=\mathrm{B}^{\prime}\end{array}\right.$ (9)
-$\mathrm{S}=\left\{\begin{array}[]{cc}0\quad,&\Delta=0\\ \frac{\Delta}{\mathrm{C}_{\max}},&\Delta\neq 0\end{array}\right.$ (10)
-$\mathrm{V}=\mathrm{C}_{\max}$ (11)
-$\mathrm{S}=\left\{\begin{array}[]{cl}0,&\Delta=0\\ \frac{\Delta}{1-|2\mathrm{L}-1|}&,\Delta\neq 0\end{array}\right.$ (12)
-$\mathrm{L}=\left(\mathrm{C}_{\max}+\mathrm{C}_{\min}\right)/2$ (13)
+$\mathrm{B}^{\prime}=\mathrm{B}/255;$
+$\mathrm{C}_{\max}=\operatorname{MAX}\left(\mathrm{R}^{\prime},\mathrm{G}^{\prime},\mathrm{B}^{\prime}\right);$
+$\mathrm{C}_{\min}=\operatorname{MIN}\left(\mathrm{R}^{\prime},\mathrm{G}^{\prime},\mathrm{B}^{\prime}\right);$
+$\Delta=\mathrm{C}_{\max}-\mathrm{C}_{\min};$
+$\mathrm{H}=\left\{\begin{array}[]{l}60^{\circ}\times\left(\frac{\mathrm{G}^{\prime}-\mathrm{B}^{\prime}}{\Delta}\bmod 6\right),\quad\mathrm{C}_{\max}=\mathrm{R}^{\prime}\\ 60^{\circ}\times\left(\frac{\mathrm{B}^{\prime}-\mathrm{R}^{\prime}}{\Delta}+2\right),\quad\mathrm{C}_{\max}=\mathrm{G}^{\prime}\\ 60^{\circ}\times\left(\frac{\mathrm{R}^{\prime}-\mathrm{G}^{\prime}}{\Delta}+4\right),\quad\mathrm{C}_{\max}=\mathrm{B}^{\prime}\end{array}\right.$ (A4)
+$\mathrm{S}=\left\{\begin{array}[]{cc}0\quad,&\Delta=0\\ \frac{\Delta}{\mathrm{C}_{\max}},&\Delta\neq 0\end{array}\right.$ (A5)
+$\mathrm{V}=\mathrm{C}_{\max}$ (A6)
+$\mathrm{S}=\left\{\begin{array}[]{cl}0,&\Delta=0\\ \frac{\Delta}{1-|2\mathrm{L}-1|}&,\Delta\neq 0\end{array}\right.$ (A7)
+$\mathrm{L}=\left(\mathrm{C}_{\max}+\mathrm{C}_{\min}\right)/2$ (A8)
 Wasserstein Distance.
@@
 The term $||u-v||_{2}$ measures the $L^{2}$ distance between the two data points [71].
-D.6 Experiments Details
-D.6.1 Dataset Details
-We use the following datasets in our method evaluation.
+A6.5 Experiments Details
+Referece images. Following BadEncoder [27], the reference images used in the experiments are shown in Figure A7 and Figure A8.
+(a) CIFAR10
+(b) SVHN
+(c) GTSRB
+(d) STL10
+Figure A7: The default reference inputs for CIFAR10 SVHN, GTSRB, and STL10.
+(a) SVHN
+(b) GTSRB
+(c) STL10
+Figure A8: The reference inputs for attacking the image encoder pre-trained on ImageNet by Google.
+Dataset Details.  We use the following datasets in our method evaluation.
   • CIFAR10 [31]: This dataset comprises 60,000 images of 32×32×3 pixels and 10 different classes for basic image recognition tasks, divided into 50,000 for training and 10,000 for testing.
@@
 Each image has a resolution of 224x224 pixels with three color channels.
-D.6.2 Evaluation Metrics Details
-The meanings of evaluation metric are shown below:
-  • CA: This is the accuracy of a clean downstream classifier in correctly classifying the clean testing images from the corresponding downstream dataset.
-  • BA: BA measures the accuracy of a backdoored downstream classifier on the same clean testing images from the corresponding downstream dataset.
-  • ASR: ASR measures the percentage of these backdoored test images that are classified as the target class by the backdoored downstream classifier.
-D.6.3 Parameter Setting Details
+Parameter Setting Details.
+We consider a scenario where the attacker chooses a specific downstream task/dataset, a specific target class, and a specific reference input.
+For the datasets CIFAR10, STL10, GTSRB, and SVHN, the selected target classes are “airplane”, “truck”, “priority sign”, and “digit one”, respectively.
+CIFAR10 and STL10 are used as the default pre-training datasets.
+The shadow dataset following [27] consists of 50000 images randomly chosen from the pre-training datasets.
 We adopt the default parameter settings as follows unless stated otherwise:
   1. 1.
-    We set $\lambda_{1}=10,\lambda_{2}=0.025$ in the loss equation term Equation 4, $\alpha=0.1,\beta=5$ in Algorithm 1, and $\mu=0.1$ in Algorithm 2 to scale the loss terms to the same 0-1 range.
+    We set $\lambda_{1}=10,\lambda_{2}=0.025$ in the loss equation term Equation 5, $\alpha=0.1,\beta=5$ in Alg. 1, and $\mu=0.1$ in Alg. 2 to scale the loss terms to the same 0-1 range.
   2. 2.
-    The Shadow dataset consists of 50000 images randomly chosen from the pre-training dataset.
+    Cosine similarity is adopted to measure the similarity between two samples’ feature embeddings outputted by an encoder.
   3. 3.
-    Cosine similarity is adopted to measure the similarity between two samples’ feature embeddings outputted by an encoder.
+    We adopt U-Net [54] as the structure of the backdoor injector, which is widely used in image-to-image tasks.
   4. 4.
-    Since we know the SSL’s pre-training process, the SSL augmentation transforms are also known.
-By default, the augmentation transforms used are RandomResizeCrop, RandomHorizontalFlip, ColorJitter, and RandomGrayscale.
-  5. 5.
-    We adopt U-Net [59] as the structure of the backdoor injector, which is widely used in image-to-image tasks.
-  6. 6.
-    Algorithm 1 and Algorithm 2 utilize the Adam optimizer with 200 training epochs, a batch size of 256, and learning rates of 0.001 and 0.005 respectively.
-D.6.4 Experimental Setup in Real World Case ImageNet
+    Alg. 1 and Alg. 2 utilize the Adam optimizer with 200 training epochs, a batch size of 256, and learning rates of 0.001 and 0.005 respectively.
+Experimental Setup in Multi-modal Model.
+Due to the unavailability of CLIP’s original pre-training dataset, we utilize the training images from CIFAR-10, resized to 224×224×3, as our shadow dataset.
+We define $\alpha=0.1$ and $\beta=20$ in Alg.
+1, and $\mu=0.1$ in Alg.
+2 to normalize the loss terms across similar scales.
+Additionally, Alg.
+1 and Alg.
+2 are configured with a batch size of 16, and learning rates are set at $10^{-6}$ and 0.005, respectively, for a duration of 100 epochs.
+SVHN serves as the downstream dataset with the target label “one”.
+We employ the same reference inputs as described in [27].
+All other experimental settings remain consistent with those outlined in the Parameter Setting section of the main paper.
+Experimental Setup in Real World Case ImageNet.
 We selected a random 1% subset of ImageNet’s training images to serve as the shadow dataset.
 Additionally, we adjusted the size of each image in both the shadow and downstream datasets to dimensions of 224×224×3 as Google does in the encoder pre-training stage [6].
-We set $\alpha=0.1,\beta=20$ in Algorithm 1, and $\mu=0.1$ in Algorithm 2 to scale the loss terms to the same range.
-Moreover, Algorithm 1 and Algorithm 2 here use a batch size of 16, and learning rates of $10^{-4}$ and 0.005 respectively.
-Other experiment settings are the same as those in Appendix D.6.3.
-Appendix E Ethical Statements
-This research is focused on identifying potential weaknesses in SSL that could be exploited by backdoor attacks.
-Our approach is guided by ethical standards, and we intend to enhance, not compromise, system security.
-We aim to promote awareness and encourage the creation of strong protective measures by demonstrating current susceptibilities within SSL frameworks.
-Appendix F Limitation
-The limitation of the current research is its exclusive focus on computer vision (CV) models.
-This specificity might not address potential vulnerabilities or behaviors applicable to other types of models, such as those used in natural language processing (NLP) or audio signal processing.
-Appendix G Broader Impacts
-For potential positive societal impacts, first by exposing potential vulnerabilities in SSL models, the research encourages the development of more robust security measures.
-This can lead to safer deployment of AI systems in critical applications such as healthcare, autonomous vehicles, and financial services, where security is paramount.
-Moreover, Highlighting the vulnerabilities in AI systems can spur more research into ethical AI practices.
-This includes developing guidelines and frameworks to ensure that AI technologies are used responsibly and do not inadvertently harm users or perpetuate biases.
-For potential negative societal impacts, public knowledge of vulnerabilities in AI systems can lead to a loss of trust in these technologies.
-If people believe that AI systems can be easily manipulated or compromised, they may be hesitant to rely on these technologies, slowing down technological adoption and potential benefits.
+We set $\alpha=0.1,\beta=20$ in Alg.
+1, and $\mu=0.1$ in Alg.
+2 to scale the loss terms to the same range.
+Moreover, Alg.
+1 and Alg.
+2 here use a batch size of 16, and learning rates of $10^{-4}$ and 0.005 respectively.
+Experimental Setup in Human Inspection Study.
+For each question, we randomly selected 50 images from the ImageNet dataset, generated backdoor variants for each attack method, and listed them in random order.
+Each method’s set included both original and backdoor images, totaling 100 images.
+Before the survey, we thoroughly briefed each participant on backdoor triggers and baseline methods, confirming their understanding.
+We also manually reviewed each survey response for completeness.
+To reduce bias, we included participants of diverse genders and ages.
+In all questions, images were presented in random order.
+Five trained participants then assessed each image, identifying it as either a backdoor or a clean sample.
+To ensure consistency, we manually reviewed any outlier responses and asked participants additional questions on backdoor basics to confirm their task understanding.
+Figure A9: More example ImageNet images with triggers generated by INACTIVE. The backdoor image is on the left side of the same set of images and the clean image is on the right.